### v2

# Token Embeddings Violate the Manifold Hypothesis

Abstract.
A full understanding of the behavior of a large language model (LLM) requires our understanding of its input token space. If this space differs from our assumptions, our understanding of and conclusions about the LLM will likely be flawed. We elucidate the structure of the token embeddings both empirically and theoretically. We present a novel statistical test assuming that the neighborhood around each token has a relatively flat and smooth structure as the null hypothesis. Failing to reject the null is uninformative, but rejecting it at a specific token $\psi$ implies an irregularity in the token subspace in a $\psi$-neighborhood, $B(\psi)$. The structure assumed in the null is a generalization of a manifold with boundary called a smooth fiber bundle (which can be split into two spatial regimes – small and large radius), so we denote our new hypothesis test as the “fiber bundle hypothesis.” Failure to reject the null hypothesis is uninformative, but rejecting it at $\psi$ indicates a statistically significant irregularity at $B(\psi)$. By running our test over several open-source LLMs, each with unique token embeddings, we find that the null is frequently rejected, and so the evidence suggests that the token subspace is not a fiber bundle and hence also not a manifold. As a consequence of our findings, when an LLM is presented with two semantically equivalent prompts, if one prompt contains a token implicated by our test, the response to that prompt will likely exhibit less stability than the other.
1. Introduction
A core area for Artificial Intelligence (AI) research is to understand the central components of modern learning systems, e.g. large language models (LLMs), assuming certain regularity conditions. Extensive research efforts have been mobilized to tackle the explainability of AI question ([1]), leveraging ideas from computer science ([2, 3, 4]), information theory ([5, 6]), and theoretical physics ([7, 8, 9]), to mathematical insights from high-dimensional probability theory ([10, 11]), and algebraic and differential topology ([12, 13, 14]) either to prove mathematical guarantees or to elucidate model behavior.
In many—if not most—AI research papers, a manifold hypothesis is tacitly assumed, that the data are concentrated near a low curvature manifold without boundary. This assumption drives the research just as an assumption of Gaussianity would drive data modeling. It is important to test this kind of assumption before data modeling. One common strategy to perform such a test is simply “plotting the data” to ascertain the empirical distribution. Plotting high-dimensional data, however, is non-trivial, but there are statistical tests to help us understand the underlying distribution (Mann-Whitney, Kolomogorov-Smirnoff, etc.). In this paper, we provide an analogous test for the regularity of token embeddings.
The first layer of most LLMs is usually comprised of the token embeddings, which are vector representations of tokens (fragments of human language, e.g. parts of words, or numbers). Thus, a good first step to explain the behavior of an LLM is to understand its token subspace where these embedding vectors live. For the remainder of this paper, the term “token”’ used in the context of the token subspace will refer to this embedding. Numerous papers (for instance, see [15, 16, 17, 18, 19, 20, 21]) have pointed to unexpected behaviors exhibited by LLMs that hinge on subtle changes in wording and text layout between apparently similar prompts, suggesting that certain semantically similar tokens may have dramatically different neighborhoods in the token subspace. These differences in neighborhoods correspond to places where the token subspace does not behave stably as it is irregular at such a token. Linguistically, these irregularities may correspond to polysemy or homonyms—tokens with multiple distinct meanings ([22]).
The presence of irregularities imply small changes in language can drive large variation in the token subspace. Moreover, if the token subspace of an LLM is, in fact, singular (not a manifold), irregularities can persist into the output of the LLM, perhaps unavoidably and potentially even because of its architecture. This is because the input and output of LLMs are a fixed finite set of tokens yet the internal architectures of the model assume a structure resembling a Riemannian manifold: a metrizable space that locally preserves inner products. Not accounting for singularities in the token subspace may confound our understanding of the LLM’s behavior.
1.1. Contributions
We present three hypothesis tests that determine for a given token $\psi$, whether its token subspace neighborhood $B(\psi)$ contains an irregularity, i.e. either has significantly high curvature or is singular. The tests are based on Theorem 1 that allows us to identify changes in subspace dimension that are inconsistent with the token subspace being:
  • 1.
    A manifold (without boundary) with a given embedded reach (low reach corresponds to high curvature and vice versa ([23])),
  • 2.
    A fibered manifold with boundary (which we call a fiber bundle for brevity) with the same embedded reach, estimated using small radii, and
  • 3.
    A fiber bundle with the same embedded reach, estimated using large radii.
Figure 1. This figure shows the application of this paper’s manifold hypothesis test to a neighborhood of the token ember in Mistral7B. (The fiber bundle tests are not shown due to space considerations.) Since the latent space of Mistral7B has dimension $\ell=4096$, we have projected the coordinates of each token shown to $3$ using principal components analysis (PCA). The inset shows all of the tokens in Mistral7B (rendered with t-SNE), with a crosshairs showing the location of the zoomed-in region. Each point shown in this figure is a token used by Mistral7B, colored based upon the $p$-value for the manifold hypothesis test. Darker colors correspond to lower $p$-values, which are less consistent with the hypothesis that the token’s neighborhood is a manifold with low curvature. Putative geometric structures are labeled for certain tokens.
When these tests are rejected, this implies that $\psi$ is less stable under mathematical transformations than other tokens. In Section 3.2, we show the endemic nature of these instabilities by proving that context cannot resolve singularities persistently. And finally, we provide conditions (Theorem 2) for when singularities can propagate into the output of an LLM for generic transformers.
In our work, we provide empirical evidence that the global token embedding obtained from four different LLMs cannot be a manifold due to non-constant local dimension. Because our hypothesis tests require the distances between all pairs of tokens, we need the token embedding vectors. For this reason, we conduct our analysis to four open source models of moderate size – GPT2 ([24]), Llemma7B ([25]), Mistral7B ([26]), and Pythia6.9B ([27]). We emphasize that while this paper demonstrates the manifold and fiber bundle tests on four open source LLMs, our main contribution is primarily a methodology for hypothesis testing on embedded points. In particular, our method can be applied to any point cloud data set; the token embeddings we study here are merely one example.
Figure 2. (Top) Flowchart of our statistical methodology. (Bottom) The distribution of local dimensions estimated near tokens in GPT2, produced by our method as an intermediate step. This histogram shows a mixture with at least three distinct peaks.
1.2. Background and related work
Manifold learning takes its roots in nonlinear dimensionality reduction of high dimensional data into low dimensional representations of smooth manifolds. While these algorithms have played an important role in modern data science especially in domains such as cosmology ([28]), astrophysics ([29, 30]), dynamical systems ([31, 32]), and quantum chemistry ([33]), the main utility of these methods have been for data visualization. For instance, even though the Uniform Manifold Approximation and Projection (UMAP) provided an optimal algorithm for fitting sampled data to a manifold, the authors themselves concede that “if the global structure is of primary interest then UMAP may not be the best choice for dimensionality reduction” ([34]). These manifold learning algorithms will find some smooth structure to fit the data rather than “error out” when the data is not sampled from a manifold as it is not a hypothesis test. While UMAP will fit a manifold to data optimally, for instance, the manifold it finds may have extremely high curvature or dimension. Both of these are undesirable properties in a model, and both implicitly suggest that a manifold may be a poor choice for a representation. To address this concern, [35] presented a test for whether data lie on a manifold with a specified embedded reach. Unfortunately, [35] does not include the values of key constants, which makes the practical usage of this test impossible. Our paper remedies both issues: we construct a practical hypothesis test for manifolds and fiber bundles with a specified embedded reach and implement it as software.
In our test case, the basic assumption is that transformers act on the entire extrinsic latent space containing the token embedding. This is clearly a manifold as it is Euclidean. But the truth is that a transformer in the context of an LLM really only acts on the embedded token subspace. The token input embedding matrix defines this subspace, by specifying where the tokens are located. Such a space is quite unconstrained. There is no a priori reason to suspect it is a manifold. It has already been shown that local neighborhoods of each token have salient topological structure ([13]), and one of the most basic parameters is the dimension near any given token in this space. Higher dimension at a token $\psi$ means that $\psi$ has greater density while lower dimensional tokens have less density ([22]). Intuitively, a densely populated $B(\psi)$ will generically occupy a higher dimension than otherwise.
Assuming that word embeddings yield manifolds, some researchers have used global dimension estimators on token input embeddings and word embeddings ([36, 37]). By using a local (not global) dimension estimator, [38] presented the first (to our knowledge) direct test of whether the token subspace is a manifold for the token input embeddings for several LLMs, though no rigorous hypothesis test was presented. Nevertheless, a strongly negative result was obtained: the subspace of tokens is apparently never a manifold, so global dimension estimators are not reliable.
2. Methods
Our method, summarized in Figures 2 and 3, proposes three related hypotheses: (1) the manifold hypothesis, (2) the fiber bundle hypothesis for small radius, and (3) the fiber bundle hypothesis for larger radius.
Figure 3. (a) Schematic of our three hypothesis tests (b) Application of the tests to the tokens #, ¢, and \$ with GPT2’s token input embedding.
The key insight is that Theorem 1 asserts that for a fiber bundle, the number of tokens within a ball of a given radius $r$ of a fixed token is a piecewise log-log-linear function of radius, in which the slopes correspond to dimension and decrease as $r$ increases. If there is a statistically significant change in the slope estimates (here, the significance level $\alpha=10^{-3}$), then we reject the manifold hypothesis. If there is a statistically significant increase in slope, then we additionally reject the fiber bundle hypothesis in the corresponding region (small or large radius).
Figure 3(b) demonstrates our method on three tokens used by GPT2. The tokens \$ and # play an important role in many programming languages, whereas ¢ does not. (Note that these tokens were chosen for illustrative purposes. The $p$-value for rejecting the hypotheses is larger than $\alpha=10^{-3}$ used in Section 3.3, so these particular tokens do not appear in any of the tables in the following sections.) The token # does not show any rejections of our hypotheses. While this does not allow one to conclude that the vicinity of # is a manifold or fiber bundle, we can use Theorem 1 to estimate the dimension from the slope of the curve (approximately $53$, once logarithms of the values on both axes are taken). The token ¢ exhibits a rejection of the manifold hypothesis, but not the fiber bundle hypothesis. We can conclude that there is a irregularity at a distance of roughly $3.9$ units of ¢. The curve for \$ exhibits two slope changes at radii $3.8$ and $4.0$. The smaller radius slope change represents a violation of the manifold hypothesis, while the large radius slope change represents a violation of the fiber bundle hypothesis.
3. Results
3.1. Volume versus radius for fiber bundles
The learned token embedding vectors in an LLM each have the same number of elements, which we refer to as $\ell$. Thus, the explicit representation of the token subspace $T$ used by an LLM arises by embedding $T$ within the latent space $\mathbb{R}^{\ell}$ by way of a smooth function $e:T\to\mathbb{R}^{\ell}$. Since embeddings are bijective, each $\psi\in T$ corresponds to a unique element $e(\psi)\in\mathbb{R}^{\ell}$. With a slight abuse of notation, we can write $\psi\in e(T)$.
The fiber bundle model imposes a specific assumption about the structure on $T$, by relating it to another, lower dimensional, manifold $S$, called the base space, via a smooth map $p:T\to S$.
Definition 1.
A fibered manifold is a surjective function $p:T\to S$ such that the Jacobian matrix $J_{\psi}p$ at every point $\psi\in T$ has rank equal to the dimension of $S$. For brevity, we refer to this kind of space simply as a fiber bundle.
If a fiber bundle is the correct representation of the space of tokens $T$, then the probability distribution of tokens given by the volume form $v$ on $T$ will impact the distribution of tokens within $\mathbb{R}^{\ell}$.
Theorem 1.
Suppose that $T$ is a compact, finite-dimensional Riemannian manifold with boundary [^1] Every point in a manifold with boundary has a neighborhood that is locally homeomorphic to a half-space. As a consequence, manifolds are a special case of manifolds with boundary., with a volume form $v$ satisfying $v(T)<\infty$, and let $p:T\to S$ be a fibered manifold. If $e:T\to\mathbb{R}^{\ell}$ is a smooth embedding with reach $\tau$, then there is a function $\rho:e(T)\to[0,\tau]$ such that if $\psi\in e(T)$, the induced volume $(e_{*}v)$ in $\mathbb{R}^{\ell}$ satisfies
(1) $(e_{*}v)\left(B_{r}(\psi\right)=\begin{cases}O\left(r^{\dim T}\right)&\text{if}0\leq r\leq\rho(\psi),\\ (e_{*}v)\left(B_{\rho(\psi)}(\psi)\right)+O\left((r-\rho(\psi))^{\dim S}\right)&\text{if }\rho(\psi)\leq r,\\ \end{cases}$
where $B_{r}(\psi)$ is the ball of radius $r$ centered at $\psi$, and the asymptotic limits are valid for small $r$.
Proof.
See Section A.4
Theorem 1 characterizes the resulting probability distribution using parameters (the exponents in Equation (1)) that can be estimated from the token input embedding. These parameters are bounded by the dimensions of $S$ and $T$.
In short, if we take the logarithm of both the volume and radius in Equation (1), we obtain a piecewise linear curve, in which the slopes are given by $\dim S<\dim T$, and the slopes must decrease through a discontinuity as the radius increases, as shown in Figure 3(a).
3.2. Context does not persistently resolve singularities
Given a context window of $w-1$ tokens, an LLM uses its learned parameters to predict the next ($w$-th) token. This gets added to the context window and the process continues. (See Section A.1 for details.) In this section, we provide a mathematical argument showing that the use of context is insufficient to resolve singularities present in the token subspace for a generic transformer in a persistent manner. Again, denote the token subspace by $T$, noting that Section 3.3 shows that $T$ is unlikely to be a manifold for typical LLMs. Let us begin with a lemma:
Lemma 1.
If $X$ is a manifold, $Y$ is an arbitrary topological space, and $X\times Y$ is a manifold, then $Y$ must also be a manifold.
Proof.
Observe that $Y$ is homeomorphic to $(X\times Y)/X$, which is the quotient of two manifolds via the projection map, which is a surjective submersion. Hence $Y$ is a manifold. ∎
Consequently, the contrapositive of Lemma 1 states that if $T$ is not a manifold but $T^{w-1}$ is a manifold, it follows that the space of sequences of tokens [^2] Caution: $T^{w}$ is the sequence of tokens embedded individually, which is different from the embedding of the entire sequence. Theorem 2 asserts conditions under which $T^{w}$ has a direct impact on LLM output. $T^{w}$ cannot be a manifold. At best, the latent space for every other context window size can be a manifold. For every proposed context window length, a longer window will exhibit singularities. In short, lengthening the context window does not persistently resolve singularities in the token embedding.
The remaining question is whether these singularities persist into the output of the LLM. Can the user elicit the instabilities caused by singularities? We provide sufficient conditions for unstable outputs:
Theorem 2.
Let $Z$ be the $d$-dimensional bounding manifold [^3] Since $T\subseteq\mathbb{R}^{\ell}$, it is always contained within such a bounding manifold $Z\subseteq\mathbb{R}^{\ell}$. for the token subspace, such that $T\subseteq Z$. Consider an LLM with a context window of size $w$, in which the latent space of tokens is $\mathbb{R}^{\ell}$, and we collect $m$ tokens as output from this LLM.
Suppose the following:
  1. (1)
    (enough tokens are collected from the response) $m>\frac{2wd}{\ell}$, but
  2. (2)
    (the context window is longer than the number of tokens we collected) $w\geq m$.
Under these conditions, a generic set of transformers yield a topological embedding of $T^{w}$ into the output of the LLM.
Proof.
See Section A.4
As a consequence, there always exists such an $m$ satisfying the hypothesis if $2d<\ell$, i.e. the bounding manifold has dimension less than half the latent space dimension. If our hypothesis tests show that the token subspace contains singularities and the hypotheses of Theorem 2 are satisfied, singularities will persist into the output of the LLM for generic transformers even for large context window sizes.
3.3. Experimental results
We applied our test to the token subspaces of the four open source LLMs introduced in Sec. 4.1: GPT2, Llemma7B, Mistral7B, and Pythia6.9B. In each LLM we tested, we found evidence that the token subspace is not globally a manifold because the local dimension is highly variable. Moreover, we can infer that they are either not fiber bundles or have high curvature locally; in either case—violating the hypotheses or having high curvature—model behavior near certain tokens will be highly unstable.
We observe highly statistically significant differences in the irregular tokens between LLMs—even for those with identical sets of tokens overall—which indicates that their respective training methodologies have a strong impact on the token subspace. Given these differences, prompts containing irregular tokens will likely produce dissimilar outputs across the four LLMs corroborating the results shown by [17]. Additionally, prompts containing irregular tokens within an LLM will likely be unstable and produce highly variable outputs with replicated queries. This is clear if $B(\psi)$ contains a singularity but will also be true for regions of high curvature resulting in instability for both interpolation and extrapolation tasks ([39]).
Several non-manifold features are visibly present in Figure 1. The tokens “ember” and “P” appear to be both putative examples of cusp points, places where the space exhibits a sharp exposed point. Manifolds cannot have cusp points, so attempting to approximate a cusp with a manifold—for instance, with UMAP—necessitates high curvature. The tokens “:” and “ly” appear to be pinch points—like what are present in a chain of sausages—which again cannot be approximated by a manifold unless it has high curvature.
Several additional views of Figure 1 are shown in Figure 6 in the Supplement. The coordinates, token text, $p$-values, and a video of the space with commentary are available in the ZIP file Supplement.
Table 1. Dimensional data for and number of tokens rejecting the manifold and fiber bundle hypotheses at two slope changes. $n$ is the number of tokens (vocabulary size) in each model. The $p$-values displayed report the minimum $p$-value for each test, accounting for multiple testing. Quartiles (Q2 is the median) are shown for the distribution of the dimensions estimated from the slopes of the log-volume versus log-radius plots aggregated over all tokens. The “rejects” columns list the number of tokens rejecting each hypothesis. Table 3 (resp. Table 49) in the Supplementary material shows each of the fiber bundle (manifold hypothesis) violations for each model that is listed here.
Model Manifold Fiber bundle
rejects Smaller Radius Larger radius
dim. rejects dim. rejects
GPT2 $68$ Q1: $20$ $12$ Q1: $8$ $7$
$n=50257$ $p\approx 3\times 10^{-8}$ Q2: $389$ $p\approx 9\times 10^{-6}$ Q2: $14$ $p\approx 3\times 10^{-8}$
Q3: $531$ Q3: $32$
Llemma7B $33$ Q1: $4096$ 00 Q1: $8$ $1$
$n=32016$ $p\approx 5\times 10^{-9}$ Q2: $4096$ N/A Q2: $11$ $p\approx 3\times 10^{-4}$
Q3: $4096$ Q3: $14$
Mistral7B $40$ Q1: $9$ $1$ Q1: $5$ $2$
$n=32016$ $p\approx 3\times 10^{-7}$ Q2: $48$ $p\approx 8\times 10^{-4}$ Q2: $6$ $p\approx 8\times 10^{-5}$
Q3: $220$ Q3: $9$
Pythia6.9B $54$ Q1: $2$ 00 Q1: $2$ 00
$n=50254$ $p\approx 2\times 10^{-7}$ Q2: $108$ N/A Q2: $5$ N/A
Q3: $235$ Q3: $145$
Table 1 shows the results of our hypothesis tests for the four models we analyzed. It is clear that the models have quite different token input embeddings, and all of them exhibit highly significant rejections of the manifold hypothesis. GPT2, Llemma7B and Mistral7B also reject the fiber bundle hypothesis. The rejections of the fiber bundle hypothesis are more frequent in the larger radius region than the smaller radius region, which is consistent with the polysemy interpretation of [22].
While most of the tokens are not shared between the LLMs, Llemma7B and Mistral7B do have identical token sets. The fact that Table 1 shows significant differences between these two models indicates that the structure of the irregularities for these these two models is quite different. This implies that their response to the same prompt is expected to be markedly different, even without considering the differences in their respective transformer stages.
Since there are many rejections of the manifold hypothesis, we summarize some general trends.
  • The GPT2 tokens at irregularities are tokens that can only appear at the beginning of words.
  • The Pythia6.9B tokens at irregularities are nearly all word fragments or short sequences of text that are quite meaningless on their own.
  • The Llemma7B and Mistral7B tokens at irregularities are a combination of the previous two: either they can only appear at the beginning of words or they are word fragments.
Table 2. Testing the hypotheses of Theorem 2: Will singularities generically persist into the output? See statement of Theorem 2 for variable definitions. “Small” and “Large” refer to the small and large radius estimates of dimensions from Table 1.
Model Latent dim Bounding dim. Context Min. output tokens Satisfied?
$\ell$ $d$ $w$ $m$ such that $w\geq m$?
Small Large Small Large Small Large
GPT2 $768$ $389$ $14$ $1024$ $1038$ $38$ No Yes
Llemma7B $4096$ $4096$ $11$ $4096$ $8193$ $23$ No Yes
Mistral7B $4096$ $48$ $6$ $4096$ $97$ $13$ Yes Yes
Pythia6.9B $4096$ $108$ $5$ $4096$ $217$ $11$ Yes Yes
Estimates of the bounding manifold dimension used in Theorem 2 for the four LLMs we studied are in the “dimension” columns of Table 1, with the most stringent being the dimension for smaller radius. Table 2 shows the results of testing the hypotheses of Theorem 2 using the medians (Q2). Mistral7B, and Pythia6.9B satisfy the hypotheses of Theorem 2 using the small radius dimension estimates, but GPT2 and Llemma7B do not. All four LLMs satisfy the hypotheses of Theorem 2 if we instead use the larger radius dimension estimates. This means that it is likely that the transformers of Mistral7B, and Pythia6.9B are unable to resolve the singularities that we show exist in the token subspace, so singularities will persist in the output of the model. If the true bounding manifold for GPT2 or Llemma7B is smaller than the small radius estimate suggests, these too will likely produce singularities in their output.
4. Discussion
None of the four LLMs we studied are manifold globally given non-constant local dimension. Locally, the token subspace cannot be manifolds with large reach, and three of the four also cannot be fiber bundles with large reach. Irregularities—tokens that cause rejections of the manifold hypothesis—occur in different ways across all four LLMs. Additionally, irregularities corresponding to violations of the fiber bundle hypothesis are tokens whose neighborhoods exhibit a dependency between the large- and small-scale variability.
The grammatical role of tokens is likely a root cause for some of sensitivity of LLMs to their prompts that has been observed in the literature ([20, 21]), and may explain why “explaining LLM behavior” remains challenging and open. Most methods for explaining LLM behavior in terms of dynamical systems, for instance, derive their inferential power from assuming that the token subspace is a manifold ([40, 41, 42]). Our results indicate that these theoretical methods may not apply to current LLMs.
One might argue that context helps to disambiguate and “fix” these problems, but this intuitive argument seems not always possible as context alone is insufficient to resolve singularities. As we showed in Section 3.2, singularities in the token subspace persistently lead to singularities in the space of token sequences, and under appropriate conditions these propagate into the output of an LLM.
Irregularities—which may in fact be singularities—are noticeable in a PCA projection of the token subspace (Figure 1 and Figure 6). Intuitively, this means that the use of inner products to perform token de-embeddings by orthogonal projections may not be accomplishing the intended objectives. Even small perturbations within the latent space may result in instabilities in these projections, providing potential explanations for the instabilities in LLM outputs.
In addition to token instability, the fact that token subspaces are not manifolds and may have small reach means that the geodesic distance between tokens can also be very unstable. While the distance along geodesics can be defined, it may not correlate with any sense of semantic distance between tokens. Furthermore, as [38] found, in most of the models, there are tokens with dimension 00 neighborhoods. These tokens are theoretically isolated, implying that the token subspace is disconnected and hence intrinsically infinitely far from other tokens.
Singularities may arise either as artifacts of the training process or from features of the languages being represented. Consistent with the idea that polysemy may yield singularities ([22]), several of the tokens where the fiber bundle hypothesis is rejected are clear homonyms. For instance, both “affect” and “monitor” can be used as either nouns or verbs, and their meanings are different in these two roles. Tokens also form fragments of text, and so, a token may correspond to homonyms after the addition of a prefix or suffix. A token like “aunder” can be prefixed to yield the word “launder”, a word with multiple meanings of opposite sense. Specifically, one can “launder” clothing (positive connotation) or “launder” money (negative connotation). Several other tokens where the fiber bundle hypothesis is violated form words with substantially different meanings or grammatical roles upon adding a prefix or suffix. For instance, “wins” can function as a noun, as a verb, or form the adjective “winsome”.
The differences in how the manifold and fiber bundle hypotheses are rejected across different LLMs suggest that the training methodology for each model leaves an indelible fingerprint (or gDNA coined by [17]). Making general assertions about LLMs without consideration of their training and evolution is likely fraught. Even between Llemma7B and Mistral7B, which have identical token sets, prompts likely cannot be “ported” from one LLM to another without significant modifications if they contain tokens near irregularities. This may have serious ramifications for interacting LLMs such as those powering multi-agentic frameworks being deployed today ([43, 44, 45]).
A few clear patterns among tokens near irregularities are nevertheless noticeable. Tokens that begin a word or are a word fragment are often located at an irregularity. Additionally, in Llemma7B (but not Mistral7B) and Pythia6.9B the tokens with unusually low fiber dimension often contain non-printing or whitespace characters. This suggests that these models are quite sensitive to text layout, perhaps to the exclusion of more semantically salient features in the text. Given our findings, future experiments can be run to explore the impact of irregular tokens on the variability of responses produced by different LLMs.
As we attempt to model natural phenomena particularly in our physical world, we know singularities exist (e.g. phase transitions), and so assuming regularity limits our understanding of the natural world. Our work extends this notion to the linguistics of natural language where its richness and complexity do not appear to manifest in a smooth representation.
4.1. Limitations
While topological theory motivates our statistical test, only the slope changes that occur for radii less than the embedded reach correspond to rejections of the manifold or fiber bundle hypotheses. If the reach of a manifold is small, then it likely has a high local curvature, and measuring distances on that surface will be numerically unstable. As a result, the reach itself is difficult to estimate accurately from sampled data ([23, 46, 47]). Our implementation tests the radius below an estimated reach, determined by an inspection of the radius-vs-volume plots of a random sample of tokens. If the true reach is smaller than our estimates, the curvature of the manifolds that fit the data well, found by any method, will be high.
And lastly, we emphasize that these tests allow us to reject the structure assumed in our null hypothesis (smooth and regular), but do not permit us to infer an alternative. While rejecting regularity does imply singularity, it does not afford us any evidence as to the type of singularity. That must be determined post hoc if possible.
Acknowledgments
The authors would like to thank Mohammed Abouzaid, Anand Sarwate, Andrew Lauziere, and Andrew Engel for helpful suggestions on a draft of this manuscript.
This material is based upon work partially supported by the Defense Advanced Research Projects Agency (DARPA) under Contract No. HR001124C0319. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the Defense Advanced Research Projects Agency (DARPA). Distribution Statement “A” (Approved for Public Release, Distribution Unlimited) applies to the portion of the work funded by DARPA.
Appendix A Supplementary Material
A.1. Background
At an abstract but precise level, an LLM consists of several interacting processes, as outlined in Figure 4. An LLM implements a transformation of a sequence of tokens (the query) into a new sequence of tokens (the response). Formally, if each input token is an element of a metric space $T$, then the LLM is a transformation $T^{w}\to T^{m}$, where $w$ is the number of tokens in the query and $m$ is the number of tokens in the response. This transformation is typically not a function because it is stochastic—it involves random draws.
Figure 4. Data flow in a typical LLM. A sequence of tokens forming the query is converted via the token input embedding $e^{w}$ into the initial context window, as a point in the latent space $X^{w}$. Each of these windows in the latent space are converted, token-by-token, into probability distributions via $f$ into the single token latent space $X$. From these, each token presented in the output (in the set $Y$) is obtained via a random draw. These output tokens are then used for subsequent windows.
To operate upon tokens using numerical models, such as could be implemented using neural networks, we must transform the finite set of tokens $T$ into numerical data. This is typically done by way of a pair of latent spaces $X=\mathbb{R}^{\ell}$ and $Y=\mathbb{R}^{q}$. The dimension $q$ of $Y$ is chosen to be equal to the number of elements in $T$, so that elements of $Y$ have the interpretation of being (unnormalized) probability distributions over $T$.
The transformation $T^{w}\to T^{m}$ is constructed in several stages.
Input tokenization:
: Each token is embedded individually via the token input embedding function $e:T\to X$. As a whole, $X^{w}$ is called a latent window.
Transformer blocks:
: The probability distribution for the next token is constructed by a continuous function $f:X^{w}\to Y$. This is usually implemented by one or more transformer blocks.
Output tokenization:
: Given the output of one of the transformer blocks $f$, one can obtain an output token in $T$ by a random draw. Specifically, if $(\psi_{1},\psi_{2},\dotsc,\psi_{w})$ is the current window in $X^{w}$, then the next token $t$ is drawn from the distribution given by $f(\psi_{1},\psi_{2},\dotsc,\psi_{w})$.
Next window prediction:
: Given that token $t$ was drawn from the distribution, the next latent window itself is constructed by a transformation $F:X^{w}\to X^{w}$, which advances the window as follows:
$F(\psi_{1},\psi_{2},\dotsc,\psi_{w}):=(\psi_{2},\dotsc,\psi_{w},t).$
Note well: the function $e^{w}:T^{w}\to X^{w}$ which transforms a sequence of individual tokens into coordinates in the latent space is merely the first stage in the process. As such, $e^{w}$ is distinct from the embedding of sequences provided by many LLM software interfaces. That sequence embedding is properly an intermediate stage within the function $f$.
The focus of this paper is specifically upon the structure of the token input embedding $e:T\to X=\mathbb{R}^{\ell}$. Since the token set $T$ is finite, $e$ can be stored as a matrix. In this matrix, each column corresponds to an element of $T$, and thereby ascribes a vector of numerical coordinates to each token. By replacing the last layer of the deep neural network $f:X^{w}\to Y$, a vector of probabilities for the next token is obtained from the activations of the last layer. One can therefore interpret the probabilities as specifying a token output embedding. Both the tokenization and the transformer stages are learned during training, and many strategies for this learning process are discussed extensively in the literature. These two stages interact when they produce the LLM output, so it is important to understand the lineage of a given tokenization as being from a particular LLM.
A.2. Computational resource requirements
Applying our entire method (starting with the token embedding matrix and ending with the $p$-values for each of the three tests at every token) to each model took approximately 12 hours of wall clock time (per model) on an Intel Core i7-3820 with 32 GB of CPU RAM and no GPU running at 3.60GHz. Almost all of the runtime was spent computing the token distance matrix. The subsequent stages, including the three hypothesis tests, shown in Figure 2 run in only a few minutes on the same machine. The distance matrix calculation was performed using scipy.spatial.distance_matrix(), which produces a dense matrix, by computing the distance between all pairs of tokens in the latent space. Because of the relatively low runtime requirements, we did not feel the need to use sparsity-enhancing techniques to accelerate the distance matrix calculation.
A.3. Interpretation of fiber bundles in terms of spatial variability
It is usual to describe measurements as exhibiting the combined effect of multiple spatial scales. One way this might arise is if we could express each measurement as being an ordered pair (signal, noise), with the short spatial scales correlating to noise and the longer spatial scales correlating to the signal. With this interpretation, if the space of all possible signals is $S$ and the space of all noise values is $V$, we could represent the space of all possible measurements as the cartesian product $E=S\times V$. This cartesian product representation can arise for noiseless data as well, as it merely captures multiple spatial scales. In what follows, we will call $S$ the base space and $V$ the fiber space.
Figure 5. Our method applied to a fiber bundle in $\mathbb{R}^{2}$. The vertical direction is the base space (long spatial scale), while the horizontal direction represents the fibers space (short spatial scale). Gray points on the right frame show estimates from a random sampling of points in the strip; the solid line shows the theoretical area versus radius curve.
Figure 5 shows an example of this situation. It consists of a $1$-dimensional base space (the long spatial scale) and $1$-dimensional fibers (the short spatial scale), which in this case forms a narrow strip in the plane. Volumes (areas, in this case) of balls of small radius scale quadratically (slope $2$ in a log-log plot), but scale asymptotically linearly (slope $1$ in a log-log plot) for large radii. The transition between these two behaviors is detectable by way of a corner in the plot. This situation is easily and robustly estimated from the data; the gray points in Figure 5 (right) are derived from a random sampling of points drawn from the strip.
To test for multiple spatial scales, we propose a signal model that is mathematically represented by a fiber bundle. In a fiber bundle, the long spatial scales are modeled by a space $S$, but all possible measurements are modeled by a function $p:E\to S$. The idea is that fibers $p^{-1}(b)$ are still cartesian products: pairs of both spatial scales, and these are all identical up to diffeomorphism. Our method relies upon a particular geometric property of fiber bundles: we can identify if the fibers are not all identical according to when the conclusion of Theorem 1 is violated.
One possible way that a fiber bundle might arise is if the underlying noise model is homooscedastic, which is to say that the dimensionality of the noise near a token does not depend on that token. In this case, the fibers (short spatial scales) mostly correspond to noise, and the base (long spatial scales) correspond mostly to signal. Nevertheless, a fiber bundle may arise with noiseless data as well; the presence of multiple spatial scales is what matters.
On the other hand, if the underlying noise model is heteroscedastic—it depends on the token in question—then the fiber bundle model should be rejected. Figure 3(a) shows a situation that is not a fiber bundle, since there is a change in the dimension of the fiber. In the upper portion of the figure, the fiber dimension is 00 while in the lower portion the fiber dimension is $1$. This is detectable by looking at the volume versus radius plots for two samples. While both samples show corners in their volume versus radius plots, Theorem 1 establishes that the slopes always decrease with increasing radius for a fiber bundle. This is violated for the sample marked $\psi_{1}$, so we conclude that the space is not a fiber bundle. On the other hand, because the sample marked $\psi_{2}$ does not exhibit this violation, it is important to note that if a sample yields data consistent with Theorem 1, we cannot conclude that the space is a fiber bundle.
A.4. Mathematical proof of Theorems 1 and 2
This section contains mathematical justification for the fiber bundle hypothesis proposed earlier in the paper and the proof of Theorems 1 and 2. The central idea is the use of a special kind of fiber bundle, namely a fibered manifold with boundary.
By the submersion theorem ([48]), if the Jacobian matrix of a fibered manifold $p:T\to S$ at every point has rank equal to the dimension of $S$, then the preimages $p^{-1}(\psi)\subseteq T$ of each point $\psi\in T$ are all diffeomorphic to each other. These preimages form the fibers discussed in the earlier sections of the paper.
As a consequence, each point $y$ in the base space $S$ has an open neighborhood $U$ where the preimage $p^{-1}(U)$ is diffeomorphic to the product $U\times p^{-1}(y)$, which is precisely the base-fiber split discussed in Section A.3. Specifically, the base dimension is simply the dimension of $S$, whereas the fiber dimension is the $(\dim T-\dim S)$.
The notion of a fibered manifold $p:T\to S$ forms the intrinsic model of the data, which is only implicit in an LLM. The tokens present in a given LLM can be thought of as a sample from a probability distribution $v$ on $T$, which can be taken to be the Riemannian volume form on $T$ normalized so that $v(T)=1$.
Definition 2.
If $e:T\to\mathbb{R}^{\ell}$ is a continuous map and $v$ is a volume form on $T$, then the pushforward is defined by
$(e_{*}v)(V):=v\left(e^{-1}(V)\right)$
for each measurable set $V$.
It is a standard fact that if $e$ is a fibered manifold or an embedding, then $e_{*}v$ is also a volume form.
Proof.
(of Theorem 1) If $e$ is assumed to be a smooth embedding [^4] Smoothness is not a strong constraint here. Any continuous function (such as ReLU) can be approximated well enough by a smooth map., the image of $e$ is a manifold of dimension $\dim T$. The pushforward of a volume form is a contravariant functor, so this means that $e_{*}v$ is the volume form for a Riemannian metric on $e(T)$. Using this Riemannian metric on $e(T)$, then [49, Thm 3.1] implies that for every $\psi\in e(T)$, if $r\ll\tau$, then
(2) $(e_{*}v)\left(B_{r}(\psi)\right)=O\left(r^{\dim T}\right).$
Since $T$ is compact, $S$ is also compact via the surjectivity of $p$. This implies that there is a maximum radius $r_{1}$ for which a ball of this radius centered on a point on $\psi\in e(T)$ is entirely contained within $e(T)$. Also by compactness of $S$, there is a minimum radius $r_{2}$ such that a ball of radius $r_{2}$ centered on a point $\psi\in e(T)$ contains a point outside of $e(T)$.
Since $e$ is assumed to be an embedding, by the tubular neighborhood theorem ([48]), it must be that $r_{2}<\tau$. Define
$\rho(\psi):=\textrm{argmax }_{r}\left\{B_{r}(\psi)\subseteq e(T)\right\},$
from which it follows that $0<r_{1}\leq\rho(\psi)\leq r_{2}<\tau$. As a result, Equation (2) holds for all $r\leq\rho(\psi)$, which is also the first case listed in Equation (1).
If $r$ is chosen such that $\rho(\psi)<r<\tau$, the volume of the ball centered on $\psi$ of radius $r$ will be less than what is given by Equation (2), namely
$(e_{*}v)\left(B_{r}(\psi)\right)<O(r^{\dim T}).$
Since $v$ is a volume form, its pushforward $(p_{*}v)$ onto $S$ is also a volume form. Moreover, via the surjectivity of $p$,
$(e_{*}v)(B_{r}(\psi))$ $=v(e^{-1}(B_{r}(\psi)))$
$\leq v(p^{-1}(p(e^{-1}(B_{r}(\psi)))))$
$\leq(p_{*}v)(p(e^{-1}(B_{r}(\psi))))$
$\leq O(r^{\dim S}).$
From this, the second case of Equation (1) follows by recentering the asymptotic series on $\rho(\psi)$. ∎
Notice that the second case in Equation (1) may be precluded since while it holds for small $r$, it may be that $\rho(\psi)$ may not be sufficiently small. As a consequence, the second case only occurs when both $r$ and $\rho(\psi)$ are sufficiently small. In the results shown in Section 3.3, both cases appear to hold frequently.
We now address Theorem 2.
Proof.
(of Theorem 2) We need to manipulate the inequalities in the hypothesis to show that they satisfy [50, Thm. 1]. Once the hypothesis is satisfied, [50, Thm. 1] establishes that a generic set of transformers will embed $T^{w}$ into the output of the LLM. Notice that [50, Thm. 1] considers the response of the LLM to a single token, though this can be easily extended by enlarging the bounding manifold dimension. With this in mind, the hypothesis to be satisfied is that
$2wd<m\ell\leq w\ell.$
The right inequality is clearly satisfied if context window is longer than number of output tokens collected, namely $w\geq m$. The left inequality is satisfied if
$2d<\frac{m}{w}\ell,$
which can be easily rearranged to the form in the hypothesis. ∎
A.5. Additional views for Figure 1
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 6. Additional rotational views of the PCA of the neighborhood of ember in Mistral7B shown in Figure 1. Each point shown is a token. Colors are $p$-values for the manifold hypothesis, with darker colors corresponding to lower likelihood that that point’s neighborhood is a manifold with low curvature.
See Fig. 6.
A.6. Table of fiber bundle rejections in all four LLMs
Table 3. Violations of the fiber bundle hypothesis
Model Token Slope change $p$-value Comment
GPT2 laughable Smaller radius $9\times 10^{-6}$ Must start a word
GPT2 nuance Smaller radius $2\times 10^{-4}$ Must start a word
GPT2 dt Smaller radius $2\times 10^{-4}$
GPT2 Mesh Smaller radius $2\times 10^{-4}$
GPT2 affect Smaller radius $3\times 10^{-4}$ Must start a word
GPT2 Thankfully Smaller radius $3\times 10^{-4}$
GPT2 swat Smaller radius $6\times 10^{-4}$ Must start a word
GPT2 Malaysian Smaller radius $6\times 10^{-4}$ Must start a word
GPT2 Palestinian Smaller radius $7\times 10^{-4}$ Must start a word
GPT2 wins Smaller radius $8\times 10^{-4}$ Must start a word
GPT2 hedon Smaller radius $9\times 10^{-4}$
GPT2 donor Smaller radius $9\times 10^{-4}$ Must start a word
GPT2 Xan Larger radius $3\times 10^{-8}$ Must start a word
GPT2 aunder Larger radius $2\times 10^{-4}$
GPT2 Dri Larger radius $2\times 10^{-4}$
GPT2 ney Larger radius $3\times 10^{-4}$
GPT2 rodu Larger radius $3\times 10^{-4}$
GPT2 Insert Larger radius $4\times 10^{-4}$
GPT2 Ying Larger radius $4\times 10^{-4}$ Must start a word
Llemma7B pax Larger radius $3\times 10^{-4}$
Mistral7B änge Smaller radius $8\times 10^{-4}$
Mistral7B H0 Larger radius $5\times 10^{-4}$
Mistral7B monitor Larger radius $8\times 10^{-5}$ Must start a word
See Table 3.
A.7. Tables of manifold rejections in all four LLMs
Table 4. Rejections of the manifold hypothesis for GPT2 (1 of 2)
Number Token $p$-value Comment
1 Xan 0.0000000280 Must start a word
2 chemist 0.0000000602
3 fry 0.00000564 Must start a word
4 laughable 0.00000807 Must start a word
5 Retro 0.0000108 Must start a word
6 laughable 0.0000156 Must start a word
7 Thankfully 0.0000339
8 intoler 0.0000514 Must start a word
9 assumes 0.0000533 Must start a word
10 handwritten 0.0000783 Must start a word
11 MY 0.0000811 Must start a word
12 nuance 0.000105 Must start a word
13 playground 0.000110 Must start a word
14 dt 0.000124
15 aunder 0.000132
16 Mesh 0.000135
17 Disney 0.000139
18 McCl 0.000155 Must start a word
19 Dri 0.000196
20 Destiny 0.000206 Must start a word
21 ney 0.000206
22 juvenile 0.000209 Must start a word
23 advent 0.000211 Must start a word
24 perpet 0.000213 Must start a word
25 otally 0.000223
26 affect 0.000226 Must start a word
27 efficient 0.000238
28 rodu 0.000242
29 Kl 0.000260 Must start a word
30 Thankfully 0.000278
31 duct 0.000284
32 frail 0.000292 Must start a word
33 alle 0.000304 Must start a word
34 Cha 0.000322 Must start a word
35 OO 0.000333
Table 5. Rejections of the manifold hypothesis for GPT2 (2 of 2)
Number Token $p$-value Comment
36 Insert 0.000336
37 Ying 0.000357 Must start a word
38 Georgia 0.000373
39 flo 0.000384
40 uddled 0.000413
41 subsequ 0.000421 Must start a word
42 lig 0.000439 Must start a word
43 Environmental 0.000456 Must start a word
44 Vladimir 0.000466 Must start a word
45 Defence 0.000510 Must start a word
46 swat 0.000558 Must start a word
47 registering 0.000564 Must start a word
48 Malaysian 0.000569 Must start a word
49 trough 0.000575 Must start a word
50 ears 0.000615 Must start a word
51 Meh 0.000637 Must start a word
52 Palestinian 0.000648 Must start a word
53 reckless 0.000654 Must start a word
54 Wax 0.000658 Must start a word
55 Dist 0.000673 Must start a word
56 Quad 0.000725 Must start a word
57 relations 0.000749
58 wins 0.000781 Must start a word
59 KER 0.000785
60 gist 0.000786 Must start a word
61 hedon 0.000834
62 MMA 0.000839 Must start a word
63 donor 0.000865 Must start a word
64 aleb 0.000873
65 Asked 0.000873
66 Isaiah 0.000940 Must start a word
67 Survey 0.000947 Must start a word
68 youngster 0.000987 Must start a word
Table 6. Rejections of the manifold hypothesis for Llemma7B
Number Token $p$-value Comment
1 clojure 0.00000000434
2 agu 0.000000846
3 ¡Cyrillic¿ 0.000000938 Must start a word
4 custom 0.00000163
5 kW 0.00000496 Must start a word
[…]
7 porque 0.0000128 Must start a word
8 Leopold 0.0000135 Must start a word
9 Rem 0.0000232 Must start a word
10 Ern 0.0000265 Must start a word
11 though 0.0000540 Must start a word
12 č 0.0000619 Must start a word
13 endl 0.0000683 Must start a word
14 corrected 0.0000738 Must start a word
15 tw 0.0000788 Must start a word
16 ères 0.000115
17 grep 0.000151
18 kind 0.000202 Must start a word
19 pax 0.000243
20 Tri 0.000250 Must start a word
21 spielte 0.000258 Must start a word
22 Heaven 0.000262 Must start a word
23 d 0.000289
24 ¡Cyrillic¿ 0.000320
25 Card 0.000335 Must start a word
26 Wall 0.000351 Must start a word
27 periodic 0.000359 Must start a word
28 ucht 0.000463
29 jeu 0.000553 Must start a word
30 slightly 0.000807 Must start a word
31 passion 0.000823 Must start a word
32 Charles 0.000833 Must start a word
33 victory 0.000906 Must start a word
Table 7. Rejections of the manifold hypothesis for Mistral7B
Number Token $p$-value Comment
1 CV 0.000000241 Must start a word
2 ember 0.00000199
3 ¡Korean character¿ 0.00000224
4 ĭ 0.00000770
5 rise 0.0000283 Must start a word
6 poured 0.0000305 Must start a word
7 aimed 0.0000364 Must start a word
8 allocate 0.0000553 Must start a word
9 UINT 0.0000684
10 monitor 0.0000767 Must start a word
11 ¡Cyrillic¿ 0.000105
[…]
13 ¡Chinese character¿ 0.000133
14 Lat 0.000139 Must start a word
15 Vel 0.000148
16 challenges 0.000178 Must start a word
17 ¡Chinese character¿ 0.000200
18 Od 0.000222 Must start a word
19 neys 0.000244
20 selon 0.000244 Must start a word
21 measures 0.000268 Must start a word
22 assumes 0.000321 Must start a word
23 Et 0.000338 Must start a word
24 Rot 0.000438
25 ¡Chinese character¿ 0.000538
26 olutely 0.000540
27 HO 0.000567
28 Solutions 0.000598 Must start a word
29 * 0.000658
30 Condition 0.000665
31 zt 0.000677
32 iced 0.000696
33 URL 0.000722
34 SEO 0.000728 Must start a word
35 weigh 0.000728 Must start a word
36 änge 0.000765
37 Vic 0.000777 Must start a word
38 apparent 0.000780 Must start a word
39 ¡Chinese character¿ 0.000896
40 ¡0xE3¿ 0.000921
Table 8. Rejections of the manifold hypothesis for Pythia6.9B (1 of 2)
Number Token $p$-value Comment
1 psychotic 0.000000161 Must start a word
2 \\]( 0.00000147
3 1990 0.00000231
4 tiene 0.00000282 Must start a word
5 ucky 0.0000134
6 Hebrew 0.0000141 Must start a word
7 mess 0.0000184
8 reasonable 0.0000300 Must start a word
9 collector 0.0000311 Must start a word
10 embodiments 0.0000428 Must start a word
11 Inhibition 0.0000593 Must start a word
12 Hers 0.0000691 Must start a word
13 odge 0.000120
14 áĢIJ 0.000144
15 ÏĢο 0.000165
16 198 0.000170 Must start a word
17 scare 0.000192 Must start a word
18 ]{}\\_[ 0.000202
19 LIMITED 0.000212 Must start a word
20 hop 0.000232 Must start a word
21 BIO 0.000246
22 raison 0.000252 Must start a word
23 inside 0.000254
24 st 0.000261
25 waving 0.000286 Must start a word
26 Austria 0.000287 Must start a word
27 ny 0.000313
Table 9. Rejections of the manifold hypothesis for Pythia6.9B (2 of 2)
Number Token $p$-value Comment
28 bags 0.000326
29 medium 0.000372
30 incent 0.000378 Must start a word
31 carboh 0.000407 Must start a word
32 secret 0.000451 Must start a word
33 ¡Unicode characters¿ 0.000468
34 ÏĢε 0.000519 Must start a word
35 120 0.000536
36 wastes 0.000565 Must start a word
37 ressor 0.000570
38 Lip 0.000597 Must start a word
39 preview 0.000650
40 redund 0.000653 Must start a word
41 ridic 0.000673 Must start a word
42 Ramirez 0.000698 Must start a word
43 omatic 0.000707
44 Gloria 0.000733 Must start a word
45 eton 0.000780
46 chairman 0.000789 Must start a word
47 Supreme 0.000859 Must start a word
48 grazing 0.000890 Must start a word
49 activation 0.000892 Must start a word
50 imming 0.000914
51 radiological 0.000931 Must start a word
52 branc 0.000941 Must start a word
53 \$\\—\\ 0.000942 Must start a word
54 DEFAULT 0.000992
See Tables 49.

### v3

@@
 Token Embeddings Violate the Manifold Hypothesis
-Abstract.
-A full understanding of the behavior of a large language model (LLM) requires our understanding of its input token space.
-If this space differs from our assumptions, our understanding of and conclusions about the LLM will likely be flawed.
+Abstract
+A full understanding of the behavior of a large language model (LLM) requires our grasp of its input token space.
+If this space differs from our assumptions, our comprehension of and conclusions about the LLM will likely be flawed.
 We elucidate the structure of the token embeddings both empirically and theoretically.
@@
 Failing to reject the null is uninformative, but rejecting it at a specific token $\psi$ implies an irregularity in the token subspace in a $\psi$-neighborhood, $B(\psi)$.
-The structure assumed in the null is a generalization of a manifold with boundary called a smooth fiber bundle (which can be split into two spatial regimes – small and large radius), so we denote our new hypothesis test as the “fiber bundle hypothesis.” Failure to reject the null hypothesis is uninformative, but rejecting it at $\psi$ indicates a statistically significant irregularity at $B(\psi)$.
-By running our test over several open-source LLMs, each with unique token embeddings, we find that the null is frequently rejected, and so the evidence suggests that the token subspace is not a fiber bundle and hence also not a manifold.
+The structure assumed in the null is a generalization of a manifold with boundary called a smooth fiber bundle (which can be split into two spatial regimes – small and large radius), so we denote our new hypothesis test as the “fiber bundle hypothesis.” By running our test over several open-source LLMs, each with unique token embeddings, we find that the null is frequently rejected, and so the evidence suggests that the token subspace is not a fiber bundle and hence also not a manifold.
 As a consequence of our findings, when an LLM is presented with two semantically equivalent prompts, if one prompt contains a token implicated by our test, the response to that prompt will likely exhibit less stability than the other.
-1. Introduction
+1 Introduction
 A core area for Artificial Intelligence (AI) research is to understand the central components of modern learning systems, e.g.
 large language models (LLMs), assuming certain regularity conditions.
-Extensive research efforts have been mobilized to tackle the explainability of AI question ([1]), leveraging ideas from computer science ([2, 3, 4]), information theory ([5, 6]), and theoretical physics ([7, 8, 9]), to mathematical insights from high-dimensional probability theory ([10, 11]), and algebraic and differential topology ([12, 13, 14]) either to prove mathematical guarantees or to elucidate model behavior.
+Extensive research efforts have been mobilized to tackle the explainability of AI question (Gunning et al.
+(2019)), leveraging ideas from computer science (Gordon (2024); Wu et al.
+(2022); Traylor et al.
+(2021)), information theory (Tan et al.
+(2024); Wang et al.
+(2024)), and theoretical physics (Chen et al.
+(2025); Geshkovski et al.
+(2023); Lu et al.
+(2019)), to mathematical insights from high-dimensional probability theory (Huang et al.
+(2024); Asher et al.
+(2023)), and algebraic and differential topology (Bradley and Vigneaux (2025); Rathore et al.
+(2023); Bradley et al.
+(2022)) either to prove mathematical guarantees or to elucidate model behavior.
 In many—if not most—AI research papers, a manifold hypothesis is tacitly assumed, that the data are concentrated near a low curvature manifold without boundary.
-This assumption drives the research just as an assumption of Gaussianity would drive data modeling.
-It is important to test this kind of assumption before data modeling.
-One common strategy to perform such a test is simply “plotting the data” to ascertain the empirical distribution.
+This assumption drives the research just as an assumption of Gaussianity drives classical approaches to data modeling.
+For instance, the Uniform Manifold Approximation and Projection (UMAP) algorithm (McInnes et al.
+(2018)) explicitly assumes that data are sampled near a low dimensional manifold and cautions the reader that the algorithm may not be reliable if this assumption is violated.
+Yet many researchers make inferences from UMAP, rarely checking that this assumption holds, and even arguing it is unnecessary.
+This argument has been routinely rejected in the literature (Jeon et al.
+(2025); Chari and Pachter (2023)), and so it is important to test this kind of assumption before data modeling.
+One common testing strategy is simply “plotting the data” to ascertain the empirical distribution.
 Plotting high-dimensional data, however, is non-trivial, but there are statistical tests to help us understand the underlying distribution (Mann-Whitney, Kolomogorov-Smirnoff, etc.).
 In this paper, we provide an analogous test for the regularity of token embeddings.
-The first layer of most LLMs is usually comprised of the token embeddings, which are vector representations of tokens (fragments of human language, e.g.
-parts of words, or numbers).
-Thus, a good first step to explain the behavior of an LLM is to understand its token subspace where these embedding vectors live.
+The first layer of most LLMs is usually comprised of the token embeddings, which are vector representations of tokens (fragments of texts).
+Thus, a good first step to explain the behavior of an LLM is to understand its initial token subspace where these embedding vectors lie.
 For the remainder of this paper, the term “token”’ used in the context of the token subspace will refer to this embedding.
-Numerous papers (for instance, see [15, 16, 17, 18, 19, 20, 21]) have pointed to unexpected behaviors exhibited by LLMs that hinge on subtle changes in wording and text layout between apparently similar prompts, suggesting that certain semantically similar tokens may have dramatically different neighborhoods in the token subspace.
-These differences in neighborhoods correspond to places where the token subspace does not behave stably as it is irregular at such a token.
-Linguistically, these irregularities may correspond to polysemy or homonyms—tokens with multiple distinct meanings ([22]).
+Numerous papers (for instance, see Verma et al.
+(2024); He et al.
+(2024); Vargas et al.
+((to appear); Li et al.
+(2024); Battle and Gollapudi (2024); Valmeekam et al.
+(2023); Sclar et al.
+(2023); Chao et al.
+(2023)) have pointed to unexpected behaviors exhibited by LLMs that hinge on subtle changes in wording and text layout between apparently similar prompts, suggesting that certain semantically similar tokens may have dramatically different topological neighborhoods in the token subspace.
+Linguistically, these irregularities may correspond to polysemy or homonyms (Jakubowski et al.
+(2020)).
 The presence of irregularities imply small changes in language can drive large variation in the token subspace.
-Moreover, if the token subspace of an LLM is, in fact, singular (not a manifold), irregularities can persist into the output of the LLM, perhaps unavoidably and potentially even because of its architecture.
+Moreover, if the token subspace of an LLM is, in fact, singular, hence not a manifold, irregularities can persist into the output of the LLM, perhaps unavoidably and potentially even because of its architecture.
 This is because the input and output of LLMs are a fixed finite set of tokens yet the internal architectures of the model assume a structure resembling a Riemannian manifold: a metrizable space that locally preserves inner products.
-Not accounting for singularities in the token subspace may confound our understanding of the LLM’s behavior.
-1.1. Contributions
-We present three hypothesis tests that determine for a given token $\psi$, whether its token subspace neighborhood $B(\psi)$ contains an irregularity, i.e.
-either has significantly high curvature or is singular.
-The tests are based on Theorem 1 that allows us to identify changes in subspace dimension that are inconsistent with the token subspace being:
-  • 1.
-    A manifold (without boundary) with a given embedded reach (low reach corresponds to high curvature and vice versa ([23])),
-  • 2.
-    A fibered manifold with boundary (which we call a fiber bundle for brevity) with the same embedded reach, estimated using small radii, and
-  • 3.
-    A fiber bundle with the same embedded reach, estimated using large radii.
-Figure 1.
-This figure shows the application of this paper’s manifold hypothesis test to a neighborhood of the token ember in Mistral7B.
-(The fiber bundle tests are not shown due to space considerations.) Since the latent space of Mistral7B has dimension $\ell=4096$, we have projected the coordinates of each token shown to $3$ using principal components analysis (PCA).
-The inset shows all of the tokens in Mistral7B (rendered with t-SNE), with a crosshairs showing the location of the zoomed-in region.
-Each point shown in this figure is a token used by Mistral7B, colored based upon the $p$-value for the manifold hypothesis test.
-Darker colors correspond to lower $p$-values, which are less consistent with the hypothesis that the token’s neighborhood is a manifold with low curvature.
-Putative geometric structures are labeled for certain tokens.
-When these tests are rejected, this implies that $\psi$ is less stable under mathematical transformations than other tokens.
+Not accounting for singularities in the token subspace may thereby confound our understanding of the LLM’s behavior.
+1.1 Contributions
+Our contributions are: (1) two novel statistical tests (for manifolds and for fiber bundles; see below), (2) evidence to reject both hypotheses for several LLMs (Section 5.2), and therefore (3) the identification of otherwise unreported structure in the embeddings for these LLMs.
+We emphasize that while this paper demonstrates the manifold and fiber bundle tests on token embeddings, our main contribution is primarily a methodology for hypothesis testing on embedded points.
+Because connected manifolds have a unique dimension, and because the dimension of fiber bundles are characterized by Theorem 1, our two tests address the following hypotheses given in Table 1.
+Table 1: Hypotheses we test. We assume the token subspace has reach $\tau>0$ (see definition in Sec. 2), and estimate dimension in the ball centered at token $\psi$ with radius $r<\tau$.
+Manifold test Fiber bundle test
+$H_{0}$ There is a unique dimension at $\psi$ The dimension at $\psi$ in a ball of radius $r$
+does not increase as $r$ increases.
+$H_{1}$ There is not a unique dimension at $\psi$ The dimension at $\psi$ increases at some $r$
+Both tests are implemented as Algorithm 1.
+The embedded reach $\tau$ defines an assumption on the curvature of the underlying space.
+Low reach corresponds to high curvature and vice versa (Berenfeld et al.
+(2022)).
+Classical Lagrange interpolation asserts that any finite sample of points can be fit exactly by a manifold.
+The resulting manifold may have extremely high curvature, which is undesirable because it results in numerical instability (Isaacson and Keller (2012)).
+To demonstrate our method, Figure 3(a)–(c) shows the application of the manifold test to several synthetic examples.
+In Section 5.2, we then provide empirical evidence that the token embedding function obtained from each of four different open source LLMs of moderate size—GPT2 (Radford et al.
+(2019)), Llemma7B (Azerbayev et al.
+(2024)), Mistral7B (Jiang et al.
+(2023)), and Pythia6.9B (Biderman et al.
+(2023))—cannot be a manifold due to non-constant local dimension as well as tokens (points) without an intrinsic dimension.
+When these tests are rejected at token $\psi$, this implies that $\psi$ is less stable under mathematical transformations than other tokens.
 In Section 3.2, we show the endemic nature of these instabilities by proving that context cannot resolve singularities persistently.
 And finally, we provide conditions (Theorem 2) for when singularities can propagate into the output of an LLM for generic transformers.
-In our work, we provide empirical evidence that the global token embedding obtained from four different LLMs cannot be a manifold due to non-constant local dimension.
-Because our hypothesis tests require the distances between all pairs of tokens, we need the token embedding vectors.
-For this reason, we conduct our analysis to four open source models of moderate size – GPT2 ([24]), Llemma7B ([25]), Mistral7B ([26]), and Pythia6.9B ([27]).
-We emphasize that while this paper demonstrates the manifold and fiber bundle tests on four open source LLMs, our main contribution is primarily a methodology for hypothesis testing on embedded points.
-In particular, our method can be applied to any point cloud data set; the token embeddings we study here are merely one example.
-Figure 2.
-(Top) Flowchart of our statistical methodology.
-(Bottom) The distribution of local dimensions estimated near tokens in GPT2, produced by our method as an intermediate step.
-This histogram shows a mixture with at least three distinct peaks.
-1.2. Background and related work
-Manifold learning takes its roots in nonlinear dimensionality reduction of high dimensional data into low dimensional representations of smooth manifolds.
-While these algorithms have played an important role in modern data science especially in domains such as cosmology ([28]), astrophysics ([29, 30]), dynamical systems ([31, 32]), and quantum chemistry ([33]), the main utility of these methods have been for data visualization.
-For instance, even though the Uniform Manifold Approximation and Projection (UMAP) provided an optimal algorithm for fitting sampled data to a manifold, the authors themselves concede that “if the global structure is of primary interest then UMAP may not be the best choice for dimensionality reduction” ([34]).
-These manifold learning algorithms will find some smooth structure to fit the data rather than “error out” when the data is not sampled from a manifold as it is not a hypothesis test.
-While UMAP will fit a manifold to data optimally, for instance, the manifold it finds may have extremely high curvature or dimension.
-Both of these are undesirable properties in a model, and both implicitly suggest that a manifold may be a poor choice for a representation.
-To address this concern, [35] presented a test for whether data lie on a manifold with a specified embedded reach.
-Unfortunately, [35] does not include the values of key constants, which makes the practical usage of this test impossible.
-Our paper remedies both issues: we construct a practical hypothesis test for manifolds and fiber bundles with a specified embedded reach and implement it as software.
-In our test case, the basic assumption is that transformers act on the entire extrinsic latent space containing the token embedding.
-This is clearly a manifold as it is Euclidean.
-But the truth is that a transformer in the context of an LLM really only acts on the embedded token subspace.
-The token input embedding matrix defines this subspace, by specifying where the tokens are located.
-Such a space is quite unconstrained.
+1.2 Background and related work
+Manifold learning takes its roots in nonlinear dimension reduction of high dimensional data into low dimensional representations of smooth manifolds.
+While these algorithms have played an important role in modern data science especially in domains such as cosmology (Park et al.
+(2024)), astrophysics (Makinen et al.
+(2024); Einig et al.
+(2023)), dynamical systems (Masuda and Kundu (2022); Parsons and Rogers (2015)), and quantum chemistry (Hemmateenejad et al.
+(2004)), the main utility of these methods have been for data visualization.
+For instance, even though UMAP provided an optimal algorithm for fitting sampled data to a manifold, the authors themselves concede that “if the global structure is of primary interest then UMAP may not be the best choice for dimensionality reduction”.
+A manifold learning algorithm will find some smooth structure to fit the data rather than “error out” when the data is not sampled from a manifold as it is not a hypothesis test.
+Grigsby et al.
+(2023) showed the dimension of ReLU-Nets are inhomogenous, and so when algorithms fit a manifold to data optimally, for instance, the found manifold may have extremely high curvature or dimension.
+Both of these are undesirable properties in a model, and both implicitly suggest that a manifold would be a poor choice for a representation (Isaacson and Keller (2012)).
+To address this concern, Fefferman et al.
+(2016) presented a test for whether data lie on a manifold with a specified embedded reach, but did not present the values of constants that are critical for practical implementation.
+More recently, Lim et al.
+(2023); Von Rohrscheidt and Rieck (2023) present tests for the manifold hypothesis, with promising results on low dimensional datasets.
+Our paper remedies both issues: we construct a practical implementation of hypothesis tests for manifolds and fiber bundles with a specified embedded reach.
+The token input embedding matrix defines the salient subspace of the latent space for an LLM, by specifying where the tokens are located.
 There is no a priori reason to suspect it is a manifold.
-It has already been shown that local neighborhoods of each token have salient topological structure ([13]), and one of the most basic parameters is the dimension near any given token in this space.
-Higher dimension at a token $\psi$ means that $\psi$ has greater density while lower dimensional tokens have less density ([22]).
-Intuitively, a densely populated $B(\psi)$ will generically occupy a higher dimension than otherwise.
-Assuming that word embeddings yield manifolds, some researchers have used global dimension estimators on token input embeddings and word embeddings ([36, 37]).
-By using a local (not global) dimension estimator, [38] presented the first (to our knowledge) direct test of whether the token subspace is a manifold for the token input embeddings for several LLMs, though no rigorous hypothesis test was presented.
-Nevertheless, a strongly negative result was obtained: the subspace of tokens is apparently never a manifold, so global dimension estimators are not reliable.
-2. Methods
-Our method, summarized in Figures 2 and 3, proposes three related hypotheses: (1) the manifold hypothesis, (2) the fiber bundle hypothesis for small radius, and (3) the fiber bundle hypothesis for larger radius.
-Figure 3. (a) Schematic of our three hypothesis tests (b) Application of the tests to the tokens #, ¢, and \$ with GPT2’s token input embedding.
-The key insight is that Theorem 1 asserts that for a fiber bundle, the number of tokens within a ball of a given radius $r$ of a fixed token is a piecewise log-log-linear function of radius, in which the slopes correspond to dimension and decrease as $r$ increases.
-If there is a statistically significant change in the slope estimates (here, the significance level $\alpha=10^{-3}$), then we reject the manifold hypothesis.
-If there is a statistically significant increase in slope, then we additionally reject the fiber bundle hypothesis in the corresponding region (small or large radius).
-Figure 3(b) demonstrates our method on three tokens used by GPT2.
-The tokens \$ and # play an important role in many programming languages, whereas ¢ does not.
-(Note that these tokens were chosen for illustrative purposes.
-The $p$-value for rejecting the hypotheses is larger than $\alpha=10^{-3}$ used in Section 3.3, so these particular tokens do not appear in any of the tables in the following sections.) The token # does not show any rejections of our hypotheses.
-While this does not allow one to conclude that the vicinity of # is a manifold or fiber bundle, we can use Theorem 1 to estimate the dimension from the slope of the curve (approximately $53$, once logarithms of the values on both axes are taken).
-The token ¢ exhibits a rejection of the manifold hypothesis, but not the fiber bundle hypothesis.
-We can conclude that there is a irregularity at a distance of roughly $3.9$ units of ¢.
-The curve for \$ exhibits two slope changes at radii $3.8$ and $4.0$.
-The smaller radius slope change represents a violation of the manifold hypothesis, while the large radius slope change represents a violation of the fiber bundle hypothesis.
-3. Results
-3.1. Volume versus radius for fiber bundles
+It has already been shown that local neighborhoods of each token have salient topological structure (Rathore et al.
+(2023)), and one of the most basic parameters is the dimension near any given token in this space.
+Higher dimension at a token $\psi$ means that $\psi$ has greater density while lower dimensional tokens have less density (Jakubowski et al.
+(2020)).
+Intuitively, a densely populated neighborhood $B(\psi)$ will generically occupy a higher dimension than otherwise.
+Figure 1: The histogram of local dimensions estimated near tokens in GPT2 (Robinson et al. (2024)). This histogram shows a mixture with at least three distinct peaks.
+Assuming that word embeddings yield manifolds, some researchers have used global dimension estimators on token input embeddings and word embeddings (Kataiwa et al.
+(2025); Gromov et al.
+(2024); Tulchinskii et al.
+(2023)).
+By using a local (not global) dimension estimator, Robinson et al.
+(2024) presented the first (to our knowledge) direct test of whether the token subspace is a manifold for the token input embeddings for several LLMs, although no rigorous hypothesis test was presented.
+The subspace of tokens has a multimodal distribution of dimension estimates (Figure 1), so unless the space is highly disconnected, it cannot be a manifold.
+2 Mathematical background
+A manifold is a space in which every point has a Euclidean neighborhood, and the space can be covered by taking finitely many of these neighborhoods at each point.
+It follows that if a manifold is connected, then the dimension of each of these Euclidean neighborhoods is the same; we call this the dimension of the manifold.
+Figure 3(a) shows the hollow sphere, which is an example of a $2$-dimensional manifold as a subset of $\mathbb{R}^{3}$.
+Given an embedding dimension $\ell$, embedded tokens lie on a subspace $X$ of Euclidean space $\mathbb{R}^{\ell}$, with the topology and geometry inherited from $\mathbb{R}^{\ell}$.
+Equivalently, one can consider an embedding function $e:X\to\mathbb{R}^{\ell}$, which is one-to-one and has a full rank matrix of partial derivatives (Jacobian matrix).
+Embedding functions therefore preserve topological structure.
+The inherited information includes a metric (the Euclidean distance) and a notion of volume.
+The volume form on a subspace is induced by the Euclidean volume.
+If the subspace is a manifold that is compatible with the inherited metric and volume form, we call the subspace a Riemannian manifold.
+For a $2$-dimensional Riemannian manifold, the volume form is the familiar notion of area.
+In general, if a space is a Riemannian manifold of dimension $d$, then the volume $v$ of a ball of radius $r$ centered at any point is given by an equation of the form Gray (1974),
+$v=Kr^{d}+\text{(correction terms)}r^{d+1},$ (1)
+where $K>0$ is a constant depending on $d$. If the correction terms are large, the manifold is said to have high curvature.
+The definition of a manifold is global: if even one point fails to have a Euclidean neighborhood, then the space cannot be a manifold, and such a point is called a singularity.
+One commonly encountered singularity is a boundary, in which a point has a neighborhood that is topologically equivalent to a half space.
+Spaces whose only singularities are boundaries are called manifolds with boundary.
+For instance, the solid ball in $\mathbb{R}^{3}$ is a manifold with boundary.
+Its set of singularities is the hollow $2$-dimensional sphere.
+Classifying the types of singularities up to topological equivalence of their neighborhoods is in general an open problem, though some types are common.
+Figure 3(b) shows a space with cusp and boundary singularities.
+Figure 3(c) shows a singularity associated with a change in dimension.
+In order to handle singularities, we also consider a generalization of a manifold called a fibered manifold with boundary (fiber bundle for short).
+This is a space in which every point has a neighborhood that can be written as a Cartesian product $B\times F$, where $B$ is a manifold and $F$ is a manifold with boundary.
+We call $B$ the base space and $F$ the fiber space.
+Curvature (the correction terms in Equation (1)) does not generalize cleanly to spaces with singularities.
+Instead, the embedded reach plays a similar role in what follows.
+The embedded reach (simply reach) for a space $X\subseteq\mathbb{R}^{\ell}$ is the largest distance $\tau\geq 0$ that a point $y\in\mathbb{R}^{\ell}$ can be from $X$, yet still have a unique closest point on $X$.
+For convex subspaces, $\tau=\infty$.
+Conversely, spaces with a cusp (like Figure 3(b)), have $\tau=0$.
+Manifolds with small reach have high curvature, and vice versa (Berenfeld et al.
+(2022)).
+3 Theoretical insights
+Theorem 1 asserts that for a fiber bundle, the number of tokens within a ball of a given radius $r$ centered at a fixed token $\psi$ is a piecewise log-log-linear function of radius, in which the slopes correspond to dimension, and the slopes cannot increase as $r$ increases.
+3.1 Volume versus radius for fiber bundles
 The learned token embedding vectors in an LLM each have the same number of elements, which we refer to as $\ell$.
-Thus, the explicit representation of the token subspace $T$ used by an LLM arises by embedding $T$ within the latent space $\mathbb{R}^{\ell}$ by way of a smooth function $e:T\to\mathbb{R}^{\ell}$.
-Since embeddings are bijective, each $\psi\in T$ corresponds to a unique element $e(\psi)\in\mathbb{R}^{\ell}$.
-With a slight abuse of notation, we can write $\psi\in e(T)$.
-The fiber bundle model imposes a specific assumption about the structure on $T$, by relating it to another, lower dimensional, manifold $S$, called the base space, via a smooth map $p:T\to S$.
-Definition 1.
-A fibered manifold is a surjective function $p:T\to S$ such that the Jacobian matrix $J_{\psi}p$ at every point $\psi\in T$ has rank equal to the dimension of $S$.
-For brevity, we refer to this kind of space simply as a fiber bundle.
-If a fiber bundle is the correct representation of the space of tokens $T$, then the probability distribution of tokens given by the volume form $v$ on $T$ will impact the distribution of tokens within $\mathbb{R}^{\ell}$.
+Thus, the explicit representation of the token subspace $T$ used by an LLM arises by an embedding function $e:T\to\mathbb{R}^{\ell}$ into the latent space $\mathbb{R}^{\ell}$.
+Since embeddings are one-to-one, each token $\psi\in T$ corresponds to a unique element $e(\psi)\in\mathbb{R}^{\ell}$ in the latent space.
+With a slight abuse of notation, we can write $\psi$ for both.
 Theorem 1.
-Suppose that $T$ is a compact, finite-dimensional Riemannian manifold with boundary [^1] Every point in a manifold with boundary has a neighborhood that is locally homeomorphic to a half-space.
-As a consequence, manifolds are a special case of manifolds with boundary., with a volume form $v$ satisfying $v(T)<\infty$, and let $p:T\to S$ be a fibered manifold.
+Suppose that $T$ is a compact, finite-dimensional Riemannian manifold with boundary, with a volume form $v$ satisfying $v(T)<\infty$, and let $p:T\to S$ be a fiber bundle.
 If $e:T\to\mathbb{R}^{\ell}$ is a smooth embedding with reach $\tau$, then there is a function $\rho:e(T)\to[0,\tau]$ such that if $\psi\in e(T)$, the induced volume $(e_{*}v)$ in $\mathbb{R}^{\ell}$ satisfies
-(1) $(e_{*}v)\left(B_{r}(\psi\right)=\begin{cases}O\left(r^{\dim T}\right)&\text{if}0\leq r\leq\rho(\psi),\\ (e_{*}v)\left(B_{\rho(\psi)}(\psi)\right)+O\left((r-\rho(\psi))^{\dim S}\right)&\text{if }\rho(\psi)\leq r,\\ \end{cases}$
+[…]
 where $B_{r}(\psi)$ is the ball of radius $r$ centered at $\psi$, and the asymptotic limits are valid for small $r$.
 Proof.
-See Section A.4
-Theorem 1 characterizes the resulting probability distribution using parameters (the exponents in Equation (1)) that can be estimated from the token input embedding.
+See Section A.5 for full details; a sketch follows. Taking the logarithm of both sides of Equation (1) yields a linear equation, in which $d$ is the slope.
+For the case of the fiber bundle, Equation (1) must be modified to handle the case where the ball of radius $r$ “sticks out” of the space. This causes a reduction in volume, hence a smaller slope. ∎
+The probability distribution of tokens in $T$ can be modeled by the volume form $v$.
+If a fiber bundle is the correct representation of the space of tokens, Theorem 1 characterizes the resulting probability distribution in $\mathbb{R}^{\ell}$ using parameters (the exponents in Equation (2)) that can be estimated from the token input embedding.
 These parameters are bounded by the dimensions of $S$ and $T$.
-In short, if we take the logarithm of both the volume and radius in Equation (1), we obtain a piecewise linear curve, in which the slopes are given by $\dim S<\dim T$, and the slopes must decrease through a discontinuity as the radius increases, as shown in Figure 3(a).
-3.2. Context does not persistently resolve singularities
-Given a context window of $w-1$ tokens, an LLM uses its learned parameters to predict the next ($w$-th) token.
-This gets added to the context window and the process continues.
-(See Section A.1 for details.) In this section, we provide a mathematical argument showing that the use of context is insufficient to resolve singularities present in the token subspace for a generic transformer in a persistent manner.
-Again, denote the token subspace by $T$, noting that Section 3.3 shows that $T$ is unlikely to be a manifold for typical LLMs.
-Let us begin with a lemma:
-Lemma 1.
-If $X$ is a manifold, $Y$ is an arbitrary topological space, and $X\times Y$ is a manifold, then $Y$ must also be a manifold.
-Proof.
-Observe that $Y$ is homeomorphic to $(X\times Y)/X$, which is the quotient of two manifolds via the projection map, which is a surjective submersion. Hence $Y$ is a manifold. ∎
-Consequently, the contrapositive of Lemma 1 states that if $T$ is not a manifold but $T^{w-1}$ is a manifold, it follows that the space of sequences of tokens [^2] Caution: $T^{w}$ is the sequence of tokens embedded individually, which is different from the embedding of the entire sequence.
-Theorem 2 asserts conditions under which $T^{w}$ has a direct impact on LLM output.
-$T^{w}$ cannot be a manifold.
-At best, the latent space for every other context window size can be a manifold.
-For every proposed context window length, a longer window will exhibit singularities.
-In short, lengthening the context window does not persistently resolve singularities in the token embedding.
-The remaining question is whether these singularities persist into the output of the LLM.
+In Algorithm 1, Monte Carlo estimates of volume are used, simply by counting tokens within a given radius.
+Although the convergence of Monte Carlo estimates is a strong assumption, it is also extremely well-established in the literature.
+Figure 2: Interpretation of the manifold and fiber bundle tests: the manifold test detects any slope changes in the log volume versus log radius curves, while the fiber bundle test only detects slope increases.
+In short, if we take the logarithm of both the volume and radius in Equation (2), we obtain a linear curve for a manifold.
+Abrupt slope changes imply that the space is not a manifold.
+For a fiber bundle, we obtain a piecewise linear curve, in which the slopes are given by $\dim S<\dim T$, and the slopes must decrease through a discontinuity as the radius increases, as shown in Figure 2.
+3.2 Context does not persistently resolve singularities
+The remaining question is whether singularities can persist into the output of the LLM.
 Can the user elicit the instabilities caused by singularities?
@@
 Theorem 2.
-Let $Z$ be the $d$-dimensional bounding manifold [^3] Since $T\subseteq\mathbb{R}^{\ell}$, it is always contained within such a bounding manifold $Z\subseteq\mathbb{R}^{\ell}$.
+Let $Z$ be a $d$-dimensional bounding manifold [^1] Since $T\subseteq\mathbb{R}^{\ell}$, it is always contained within such a bounding manifold $Z\subseteq\mathbb{R}^{\ell}$.
 for the token subspace, such that $T\subseteq Z$.
 Consider an LLM with a context window of size $w$, in which the latent space of tokens is $\mathbb{R}^{\ell}$, and we collect $m$ tokens as output from this LLM.
-Suppose the following:
-  1. (1)
-    (enough tokens are collected from the response) $m>\frac{2wd}{\ell}$, but
-  2. (2)
-    (the context window is longer than the number of tokens we collected) $w\geq m$.
-Under these conditions, a generic set of transformers yield a topological embedding of $T^{w}$ into the output of the LLM.
+Suppose the following, (enough tokens are collected from the response) $m>\frac{2wd}{\ell}$, but (the context window is longer than the number of tokens we collected) $w\geq m$.
+Under these conditions, a generic set of transformers yields a topological embedding of $T^{w}=T\times\dotsb T$ into the output of the LLM.
 Proof.
-See Section A.4
+See Section A.5. ∎
 As a consequence, there always exists such an $m$ satisfying the hypothesis if $2d<\ell$, i.e.
@@
 If our hypothesis tests show that the token subspace contains singularities and the hypotheses of Theorem 2 are satisfied, singularities will persist into the output of the LLM for generic transformers even for large context window sizes.
-3.3. Experimental results
-We applied our test to the token subspaces of the four open source LLMs introduced in Sec.
-4.1: GPT2, Llemma7B, Mistral7B, and Pythia6.9B.
-In each LLM we tested, we found evidence that the token subspace is not globally a manifold because the local dimension is highly variable.
-Moreover, we can infer that they are either not fiber bundles or have high curvature locally; in either case—violating the hypotheses or having high curvature—model behavior near certain tokens will be highly unstable.
-We observe highly statistically significant differences in the irregular tokens between LLMs—even for those with identical sets of tokens overall—which indicates that their respective training methodologies have a strong impact on the token subspace.
-Given these differences, prompts containing irregular tokens will likely produce dissimilar outputs across the four LLMs corroborating the results shown by [17].
-Additionally, prompts containing irregular tokens within an LLM will likely be unstable and produce highly variable outputs with replicated queries.
-This is clear if $B(\psi)$ contains a singularity but will also be true for regions of high curvature resulting in instability for both interpolation and extrapolation tasks ([39]).
-Several non-manifold features are visibly present in Figure 1.
-The tokens “ember” and “P” appear to be both putative examples of cusp points, places where the space exhibits a sharp exposed point.
-Manifolds cannot have cusp points, so attempting to approximate a cusp with a manifold—for instance, with UMAP—necessitates high curvature.
-The tokens “:” and “ly” appear to be pinch points—like what are present in a chain of sausages—which again cannot be approximated by a manifold unless it has high curvature.
-Several additional views of Figure 1 are shown in Figure 6 in the Supplement.
-The coordinates, token text, $p$-values, and a video of the space with commentary are available in the ZIP file Supplement.
-Table 1.
-Dimensional data for and number of tokens rejecting the manifold and fiber bundle hypotheses at two slope changes.
+4 Methods
+The tests we propose are implemented via Algorithm 1.
+The coordinates of the points (token embedding vectors, in Section 5.2) are fixed when we analyze the data, and they do not change from run to run.
+Algorithm 1 is fully deterministic, so there is no requirement of multiple runs.
+Centered at a given point, the sorted list of distances to all other points contains the radii at which the Monte Carlo estimate of the volume of a ball increases.
+The corresponding volumes are simply the indices of this list.
+Both tests estimate dimension as slope in the log-log progression of the entries versus the indices ($3$-point centered difference estimates via numpy.gradient()), and detect slope changes.
+Algorithm 1 Manifold and fiber bundle tests
+[…]
+[…]
+3:$W$: sliding window size
+[…]
+[…]
+[…]
+7:Set of dimension estimates
+8:procedure ManifoldAndFiberBundleTest($x_{\bullet}$, $v_{min}$, $v_{max}$, $W$)
+9:  Compute $n\times n$ pairwise distance matrix $D$ between all tokens
+10:  for Each column of $D$ do $\triangleright$ Columns correspond to token indices
+11:   Sort the column $\triangleright$ Now row indices of distance matrix are volumes, entries are radii
+12:   Retain rows $v_{min}$ through $v_{max}$
+13:   Compute log-log slopes (= dimension estimates) along the column
+14:   Run two sample $T$-test along adjacent sliding windows of size $W$ with level $\alpha$:
+Manifold test:
+Append to $p_{1}$: the $p$-value for the hypothesis that the slope is constant
+Fiber bundle test:
+Append to $p_{2}$: the $p$-value for the hypothesis that the slope decreases with row index
+15:   Store both $p$ values and slope with corresponding token (column index)
+16:  end for
+17:  Apply Holm-Bonferroni multiple test correction to both sets of $p$-values
+18:end procedure
+5 Results
+To demonstrate the validity of our tests in practice, we demonstrate them on three synthetic datasets in Section 5.1 shown in Figure 3(a)–(c).
+We then turn to the main thrust of the paper, testing LLM token embeddings in Section 5.2.
+Source code is available at Robinson (2025).
+5.1 Synthetic datasets
+Figure 3: This figure shows the application of this paper’s manifold hypothesis test on three synthetic datasets: (a) a manifold, the unit sphere in $\mathbb{R}^{3}$, (b) a fiber bundle, the cusp surface $z^{8}+x^{2}+y^{2}=0$ in $\mathbb{R}^{3}$, (c) neither, a “lollipop” space, and (d) a neighborhood of the token ember in Mistral7B.
+Points are colored by the $p$ values of the manifold hypothesis test; darker colors mean reject the manifold hypothesis.
+Figure 3(a) shows the result of applying our manifold hypothesis test to the unit sphere in $\mathbb{R}^{3}$.
+Since the sphere is a manifold, the $p$ value for every point tested is close to $1$, and no rejections of either hypothesis occur.
+Figure 3(b) shows the result of applying the manifold test to a portion of the surface given by
+$z^{8}+x^{2}+y^{2}=0,\text{ where }x^{2}+y^{2}\leq 1.$
+This surface is a fiber bundle in which the base space is simply the origin, and the fiber spaces are the radial line segments tracing out from the origin to the edge of the disk.
+It is a manifold with boundary (not a manifold), and the origin $x=y=z=0$ is a cusp singularity.
+The number of points participating in the cusp and boundary is small ($41$ points) compared with the total number of points in the dataset ($1200$ points).
+The manifold hypothesis test strongly detects the cusp singularity in Figure 3(b), with $p<5\times 10^{-8}$.
+The manifold test fails to detect the boundary ($p>0.001$).
+While using a larger $v_{max}$ in Algorithm 1 improves the test’s sensitivity to the point where the boundary is detected, it also reduces the specificity of the test for identifying which points are singularities.
+On the other hand, the fiber bundle test is not rejected at all, which is consistent with the space being a fiber bundle.
+Finally, Figure 3(c) shows a sample of $3000$ points drawn from a space shaped like a candy “lollipop,” constructed by attaching a line segment (the “stick”) to a point on the surface of a hollow sphere (the “candy”).
+The lollipop is neither a manifold nor a fiber bundle.
+On the interior of the stick, the space is a $1$-dimensional manifold.
+The candy is a $2$-dimensional manifold, except where the stick attaches.
+The entire space has exactly two singular points.
+At the very end of the stick is a singular point (a boundary point).
+Where the stick attaches to the candy is another singularity; it is a transition from the $1$-dimensional stick to the $2$-dimensional candy.
+Running the manifold hypothesis test on the lollipop reveals low $p$-values on points near the boundary of the stick ($75$ points) and at the transition ($58$ points) (both have $p<10^{-5}$).
+Due to random sampling, the exact points of singularity are not actually present in the dataset; these detections are near the two singularities.
+None of the points near the boundary point of the stick cause the fiber bundle hypothesis to be rejected with $p<10^{-5}$, while $2$ points near the transition cause the fiber bundle hypothesis to be rejected with $p<10^{-5}$.
+Since even one rejection is sufficient, the tests correctly identify that the lollipop is neither a manifold nor a fiber bundle.
+5.2 LLM token embeddings
+We applied our Algorithm 1 to the token subspaces of the four open source LLMs introduced in Sec.
+1.1: GPT2, Llemma7B, Mistral7B, and Pythia6.9B.
+In each LLM we tested, Table 2 exhibits evidence that the token subspace is not globally a manifold because the local dimension is highly variable (“dim.” columns).
+Moreover, we can infer that they are either not fiber bundles or have high curvature locally; in either case—violating the hypotheses or having high curvature—model behavior near certain tokens will be highly unstable (low $p$-values in the “rejects” columns).
+Tables 510 in the Supplementary material list every violation we found.
+Additionally, Figure 3(d) shows the application of the manifold test to the tokens of Mistral7B [^2] More views of Figure 3(d) are shown in Figure 8 in the Supplement.
+in the vicinity of the token ember.
+The manifold hypothesis is rejected at various tokens (low $p$-value, shown in darker color).
+For each LLM we studied, we chose two pairs of $v_{min}$ and $v_{max}$ parameters in Algorithm 1: one yielding a small radius neighborhood and one yielding a larger radius neighborhood.
+These parameters were chosen by inspecting a small simple random sample of tokens, with the aim of identifying where changes in dimension were likely to occur.
+See Section A.6 for details.
+This configuration results in three tests: a single test for the manifold hypothesis, the fiber bundle test for small radius, and the fiber bundle test for large radius.
+Table 2: Dimensional data and number of tokens rejecting the manifold and fiber bundle hypotheses at two slope changes.
 $n$ is the number of tokens (vocabulary size) in each model.
@@
 The “rejects” columns list the number of tokens rejecting each hypothesis.
-Table 3 (resp.
-Table 49) in the Supplementary material shows each of the fiber bundle (manifold hypothesis) violations for each model that is listed here.
-Model Manifold Fiber bundle
-rejects Smaller Radius Larger radius
+Tables 410 in the Supplementary material show each of the manifold and fiber bundle hypothesis violations we found.
+Model Manifold rejects Fiber bundle
+Smaller Radius Larger radius
 dim. rejects dim. rejects
-GPT2 $68$ Q1: $20$ $12$ Q1: $8$ $7$
-$n=50257$ $p\approx 3\times 10^{-8}$ Q2: $389$ $p\approx 9\times 10^{-6}$ Q2: $14$ $p\approx 3\times 10^{-8}$
+GPT2 $n=50257$ $66$ $p\approx 3\times 10^{-8}$ Q1: 2020 $12$ $p\approx 9\times 10^{-6}$ Q1: 88 77 $p\approx 3\times 10^{-8}$
+Q2: $389$ Q2: $14$
 Q3: $531$ Q3: $32$
-Llemma7B $33$ Q1: $4096$ 00 Q1: $8$ $1$
-$n=32016$ $p\approx 5\times 10^{-9}$ Q2: $4096$ N/A Q2: $11$ $p\approx 3\times 10^{-4}$
+Llemma7B $n=32016$ $33$ $p\approx 5\times 10^{-9}$ Q1: $4096$ 0 N/A Q1: 88 11 $p\approx 3\times 10^{-4}$
+Q2: $4096$ Q2: $11$
 Q3: $4096$ Q3: $14$
-Mistral7B $40$ Q1: $9$ $1$ Q1: $5$ $2$
-$n=32016$ $p\approx 3\times 10^{-7}$ Q2: $48$ $p\approx 8\times 10^{-4}$ Q2: $6$ $p\approx 8\times 10^{-5}$
-Q3: $220$ Q3: $9$
-Pythia6.9B $54$ Q1: $2$ 00 Q1: $2$ 00
-$n=50254$ $p\approx 2\times 10^{-7}$ Q2: $108$ N/A Q2: $5$ N/A
+Mistral7B $n=32016$ $40$ $p\approx 3\times 10^{-7}$ Q1: 99 11 $p\approx 8\times 10^{-4}$ Q1: 55 22 $p\approx 8\times 10^{-5}$
+Q2: $48$ Q2: 66
+Q3: $220$ Q3: 99
+Pythia6.9B $n=50254$ $54$ $p\approx 2\times 10^{-7}$ Q1: 22 0 N/A Q1: 22 0 N/A
+Q2: $108$ Q2: 55
 Q3: $235$ Q3: $145$
 Table 2 shows the results of our hypothesis tests for the four models we analyzed.
-It is clear that the models have quite different token input embeddings, and all of them exhibit highly significant rejections of the manifold hypothesis.
+The models consequently have token input embeddings with vastly different topology, and all of them exhibit highly significant rejections of the manifold hypothesis.
 GPT2, Llemma7B and Mistral7B also reject the fiber bundle hypothesis.
-The rejections of the fiber bundle hypothesis are more frequent in the larger radius region than the smaller radius region, which is consistent with the polysemy interpretation of [22].
-While most of the tokens are not shared between the LLMs, Llemma7B and Mistral7B do have identical token sets.
-The fact that Table 1 shows significant differences between these two models indicates that the structure of the irregularities for these these two models is quite different.
-This implies that their response to the same prompt is expected to be markedly different, even without considering the differences in their respective transformer stages.
-Since there are many rejections of the manifold hypothesis, we summarize some general trends.
+The rejections of the fiber bundle hypothesis are more frequent in the larger radius region than the smaller radius region, which is consistent with the polysemy interpretation of Jakubowski et al.
+(2020).
+The “rejects” columns of Table 2 give the number of tokens that fail our test (in the top rows of each cell).
+The listed $p$ is the overall $p$-value after Bonferroni correction.
+The “dim.” column shows the quartiles for the estimates of dimensions, at least for the tokens for which a dimension can reliably be ascribed.
+Note for each of the tokens listed in Tables 510, no dimension estimate is possible.
+Since there are many rejections of the manifold hypothesis, we summarize some general trends, obtained by manual inspection of Tables 510 in the Supplementary material, to which we direct the reader.
 The GPT2 tokens at irregularities are tokens that can only appear at the beginning of words.
@@
 The Llemma7B and Mistral7B tokens at irregularities are a combination of the previous two: either they can only appear at the beginning of words or they are word fragments.
-Table 2.
-Testing the hypotheses of Theorem 2: Will singularities generically persist into the output?
+While most of the tokens are not shared between the LLMs, Llemma7B and Mistral7B have identical token sets.
+The fact that Table 2 shows significant differences between these two models indicates that the structure of the irregularities for these two models is quite different.
+This alone implies that their response to the same prompt is expected to be markedly different, even without considering the differences in their respective transformer stages, corroborating the results shown by Vargas et al.
+((to appear).
+Table 3: Testing the hypotheses of Theorem 2: Will singularities generically persist into the output?
 See statement of Theorem 2 for variable definitions.
@@
 Llemma7B $4096$ $4096$ $11$ $4096$ $8193$ $23$ No Yes
-Mistral7B $4096$ $48$ $6$ $4096$ $97$ $13$ Yes Yes
-Pythia6.9B $4096$ $108$ $5$ $4096$ $217$ $11$ Yes Yes
+Mistral7B $4096$ $48$ 66 $4096$ $97$ $13$ Yes Yes
+Pythia6.9B $4096$ $108$ 55 $4096$ $217$ $11$ Yes Yes
 Estimates of the bounding manifold dimension used in Theorem 2 for the four LLMs we studied are in the “dimension” columns of Table 2, with the most stringent being the dimension for smaller radius.
 Table 3 shows the results of testing the hypotheses of Theorem 2 using the medians (Q2).
-Mistral7B, and Pythia6.9B satisfy the hypotheses of Theorem 2 using the small radius dimension estimates, but GPT2 and Llemma7B do not.
+Mistral7B and Pythia6.9B satisfy the hypotheses of Theorem 2 using the small radius dimension estimates, but GPT2 and Llemma7B do not.
 All four LLMs satisfy the hypotheses of Theorem 2 if we instead use the larger radius dimension estimates.
-This means that it is likely that the transformers of Mistral7B, and Pythia6.9B are unable to resolve the singularities that we show exist in the token subspace, so singularities will persist in the output of the model.
+This means that it is likely that the transformers of Mistral7B and Pythia6.9B are unable to resolve the singularities that we show exist in the token subspace, so singularities will persist in the output of the model.
 If the true bounding manifold for GPT2 or Llemma7B is smaller than the small radius estimate suggests, these too will likely produce singularities in their output.
-4. Discussion
-None of the four LLMs we studied are manifold globally given non-constant local dimension.
-Locally, the token subspace cannot be manifolds with large reach, and three of the four also cannot be fiber bundles with large reach.
-Irregularities—tokens that cause rejections of the manifold hypothesis—occur in different ways across all four LLMs.
-Additionally, irregularities corresponding to violations of the fiber bundle hypothesis are tokens whose neighborhoods exhibit a dependency between the large- and small-scale variability.
-The grammatical role of tokens is likely a root cause for some of sensitivity of LLMs to their prompts that has been observed in the literature ([20, 21]), and may explain why “explaining LLM behavior” remains challenging and open.
-Most methods for explaining LLM behavior in terms of dynamical systems, for instance, derive their inferential power from assuming that the token subspace is a manifold ([40, 41, 42]).
-Our results indicate that these theoretical methods may not apply to current LLMs.
-One might argue that context helps to disambiguate and “fix” these problems, but this intuitive argument seems not always possible as context alone is insufficient to resolve singularities.
-As we showed in Section 3.2, singularities in the token subspace persistently lead to singularities in the space of token sequences, and under appropriate conditions these propagate into the output of an LLM.
-Irregularities—which may in fact be singularities—are noticeable in a PCA projection of the token subspace (Figure 1 and Figure 6).
+6 Discussion
+The large distribution of local dimension (Figure 1 and Table 2) rejects that the token embedding is a globally smooth manifold for each of the LLMs we studied.
+Beyond non-constant dimension, we also uncovered token neighborhoods without a well-defined intrinsic dimension, a structure not previously reported in these embeddings.
+These irregular structures suggest that the topology of the embeddings are more complicated than generally assumed.
+Consequently, all four token subspaces cannot locally be manifolds with large reach, and three cannot be fiber bundles with large reach.
+Given the aggregation of training data across domains (and even across natural languages), the potential for internal contradictions (contronyms) and confounders (natural language, tweets, source code), it is not surprising to see that our hypotheses are rejected.
+The semantic role of tokens likely contributes to LLM sensitivity of prompts that has been observed in the literature (Schulz et al.
+(2025); Cui et al.
+(2025); Zhao et al.
+(2025a); Zilberman (2025); Sclar et al.
+(2023); Chao et al.
+(2023)) and may underscore why elucidating LLM behavior remains challenging and open.
+The manifold hypothesis gave both an intuitive and mathematically grounded explanation for the behaviors of learning machines (Liu et al.
+(2024); Brahma et al.
+(2015); Narayanan and Mitter (2010)) yet the instability of LLM outputs strongly suggested that these models were different.
+Our work gives conclusive empirical evidence against the manifold hypothesis for token embeddings and may offer novel ways to interpret LLM behavior as we are not in a regular setting.
+Such irregularities are noticeable in a principal components analysis (PCA) projection of the token subspace (Figures 3(d), 7 and 8) but are extremely hard to interpret a priori.
+Mathematicians are still classifying singularities in $\mathbb{R}^{3}$(Zhu et al.
+(2024)) implying that the classification problem is far from complete (even in $\mathbb{R}^{3}$).
 Intuitively, this means that the use of inner products to perform token de-embeddings by orthogonal projections may not be accomplishing the intended objectives.
@@
 While the distance along geodesics can be defined, it may not correlate with any sense of semantic distance between tokens.
-Furthermore, as [38] found, in most of the models, there are tokens with dimension 00 neighborhoods.
+Furthermore, as Robinson et al.
+(2024) found, in most of the models, there are tokens with dimension 0 neighborhoods.
 These tokens are theoretically isolated, implying that the token subspace is disconnected and hence intrinsically infinitely far from other tokens.
 Singularities may arise either as artifacts of the training process or from features of the languages being represented.
-Consistent with the idea that polysemy may yield singularities ([22]), several of the tokens where the fiber bundle hypothesis is rejected are clear homonyms.
+Consistent with the idea that polysemy may yield singularities (Jakubowski et al.
+(2020)), several of the tokens where the fiber bundle hypothesis is rejected are clear homonyms.
 For instance, both “affect” and “monitor” can be used as either nouns or verbs, and their meanings are different in these two roles.
-Tokens also form fragments of text, and so, a token may correspond to homonyms after the addition of a prefix or suffix.
+A token may also correspond to a pair of homonyms after the addition of a prefix or suffix.
 A token like “aunder” can be prefixed to yield the word “launder”, a word with multiple meanings of opposite sense.
@@
 For instance, “wins” can function as a noun, as a verb, or form the adjective “winsome”.
-The differences in how the manifold and fiber bundle hypotheses are rejected across different LLMs suggest that the training methodology for each model leaves an indelible fingerprint (or gDNA coined by [17]).
+The differences in how the manifold and fiber bundle hypotheses are rejected across different LLMs suggest that the training methodology for each model leaves an indelible fingerprint (or gDNA coined by Vargas et al.
+((to appear)).
 Making general assertions about LLMs without consideration of their training and evolution is likely fraught.
-Even between Llemma7B and Mistral7B, which have identical token sets, prompts likely cannot be “ported” from one LLM to another without significant modifications if they contain tokens near irregularities.
-This may have serious ramifications for interacting LLMs such as those powering multi-agentic frameworks being deployed today ([43, 44, 45]).
-A few clear patterns among tokens near irregularities are nevertheless noticeable.
+Even between Llemma7B and Mistral7B, which have identical token sets, prompts likely cannot be “ported” from one LLM to another without significant modifications if they contain tokens near irregularities, likely a consequence of the invariance of domain theorem.
+This may have serious ramifications for interacting LLMs such as those powering multi-agentic frameworks being deployed today (Gu et al.
+(2025); Jin et al.
+(2025); Qian et al.
+(2025)).
 Tokens that begin a word or are a word fragment are often located at an irregularity.
-Additionally, in Llemma7B (but not Mistral7B) and Pythia6.9B the tokens with unusually low fiber dimension often contain non-printing or whitespace characters.
+Additionally, in Llemma7B (but not Mistral7B) and Pythia6.9B, the tokens with unusually low fiber dimension often contain non-printing or whitespace characters.
 This suggests that these models are quite sensitive to text layout, perhaps to the exclusion of more semantically salient features in the text.
 Given our findings, future experiments can be run to explore the impact of irregular tokens on the variability of responses produced by different LLMs.
+A natural question is if we can find a token embedding that is everywhere regular.
+One way to approach an answer is to view the number of tokens in a vocabulary as a hyperparameter.
+There might be a certain vocabulary size that minimizes the number of singularities given the corpus, architecture, etc.
+Even the best vocabulary might lead to a space that could still be singular, for reasons inspired by Jakubowski et al.
+(2020).
+Consider the * symbol, which happens to be singular for Mistral7B (See Table 8, number 29).
+In natural language, it has a very distinctive meaning in terms of calling out footnotes, while in other contexts it can be used for multiplication or marking text as bold (in Markdown).
+Therefore, this character connects these different semantics.
+Suppose that each language were in fact well-described by a manifold.
+Then these manifolds would be joined together in the entire latent space at the * (and likely other points).
+As a result, this token with high likelihood would be a singularity because the dimensions for different languages are almost surely different (Tulchinskii et al.
+(2023)).
+In reality, tokens are neither in 1:1 correspondence with words nor with grammatical symbols.
+Not only do parts of words map onto a single token, but tokenization is a many to one map across the board.
+How well context / attention / transformers can unfurl these mappings is a big part of the mystery of how these models work.
+Identifying and analyzing the irregularities may help us understand the approximate limits of attention/transformers architectures.
 As we attempt to model natural phenomena particularly in our physical world, we know singularities exist (e.g.
@@
 Our work extends this notion to the linguistics of natural language where its richness and complexity do not appear to manifest in a smooth representation.
-4.1. Limitations
+7 Limitations
 While topological theory motivates our statistical test, only the slope changes that occur for radii less than the embedded reach correspond to rejections of the manifold or fiber bundle hypotheses.
 If the reach of a manifold is small, then it likely has a high local curvature, and measuring distances on that surface will be numerically unstable.
-As a result, the reach itself is difficult to estimate accurately from sampled data ([23, 46, 47]).
-Our implementation tests the radius below an estimated reach, determined by an inspection of the radius-vs-volume plots of a random sample of tokens.
-If the true reach is smaller than our estimates, the curvature of the manifolds that fit the data well, found by any method, will be high.
 And lastly, we emphasize that these tests allow us to reject the structure assumed in our null hypothesis (smooth and regular), but do not permit us to infer an alternative.
@@
 That must be determined post hoc if possible.
-Acknowledgments
+Acknowledgments and Funding Disclosure
 The authors would like to thank Mohammed Abouzaid, Anand Sarwate, Andrew Lauziere, and Andrew Engel for helpful suggestions on a draft of this manuscript.
@@
 Distribution Statement “A” (Approved for Public Release, Distribution Unlimited) applies to the portion of the work funded by DARPA.
+The authors have no competing interests to declare.
 Appendix A Supplementary Material
-A.1. Background
+A.1 Additional discussion of the literature
+Although there a number of mathematically distinct definitions of dimension, they coincide for manifolds.
+Intrinsic dimension of a point cloud has been widely studied in the literature Schweinhart [2020].
+All of the papers that address practical estimation of intrinsic dimension from data do so under the assumption that the data lie on a manifold.
+For instance, Levina and Bickel [2004] provides maximum likelihood estimates, operating under the assumption that data “can be efficiently summarized in a space of a much lower dimension, such as a nonlinear manifold.” Moreover, Facco et al.
+[2017] provides a surprisingly effective approach using extremely small neighborhoods, a fact which underlies our use of our test on a “small radius neighborhood” in Section 5.2.
+Even though we ultimately reject the manifold hypothesis, the token subspace has been studied under the assumption that it is a manifold.
+For instance Lee et al.
+[2025] observe that,“Language models often construct similar local representations as constructed from local linear embeddings.” We note that most of the tokens in a typical vocabulary fail to reject the manifold hypothesis, which means that without the deeper analysis we pursue, it is easy to overlook important differences.
+Furthermore, our Table 2 shows significant topological differences in the local representations between models.
+As introduced in Section 1.2, we mentioned that Lee et al.
+[2025] show a few things that are consistent with our findings.
+They note that “Token embeddings exhibit low intrinsic dimensions,” which agrees with our in Table 2, at least in the large radius neighborhood.
+Their estimates of dimension for GPT2 is within the confidence interval we computed.
+They do not recognize that there might be a fiber bundle structure in play, so Lee et al.
+[2025] does not estimate dimensions from a smaller radius neighborhood.
+Relevant to Theorem 2, there is a possibility that the implicit space of possible activations can be inferred from the behavior of the model Zhao et al.
+[2025b].
+Since Theorem 2 represents the context window as a product of copies of the latent space, a projection of this implicit structure is the token subspace.
+In this vein, Zhao et al.
+[2025b] is not inconsistent with other findings within the literature, for instance Robinson et al.
+[2024].
+This latter paper shows that the token subspace can be inferred (up to topological equivalence) from the responses of the LLM to single-token prompts.
+Taken together, Zhao et al.
+[2025b], Robinson et al.
+[2024] and the present one have pretty significant implications.
+Even though the token subspace may have about 0.1% of tokens that are singular, we suspect that a user will be encountering singular tokens on a regular basis when using LLMs (which seems empirically plausible).
+Estimating the rate of occurrence of a singular token remains an open question.
+A.2 LLM Background
 At an abstract but precise level, an LLM consists of several interacting processes, as outlined in Figure 4.
@@
 This transformation is typically not a function because it is stochastic—it involves random draws.
-Figure 4.
-Data flow in a typical LLM.
+Figure 4: Data flow in a typical LLM.
 A sequence of tokens forming the query is converted via the token input embedding $e^{w}$ into the initial context window, as a point in the latent space $X^{w}$.
@@
 The transformation $T^{w}\to T^{m}$ is constructed in several stages.
-Input tokenization:
+Input tokenization
 : Each token is embedded individually via the token input embedding function $e:T\to X$. As a whole, $X^{w}$ is called a latent window.
-Transformer blocks:
+Transformer blocks
 : The probability distribution for the next token is constructed by a continuous function $f:X^{w}\to Y$. This is usually implemented by one or more transformer blocks.
-Output tokenization:
+Output tokenization
 : Given the output of one of the transformer blocks $f$, one can obtain an output token in $T$ by a random draw.
 Specifically, if $(\psi_{1},\psi_{2},\dotsc,\psi_{w})$ is the current window in $X^{w}$, then the next token $t$ is drawn from the distribution given by $f(\psi_{1},\psi_{2},\dotsc,\psi_{w})$.
-Next window prediction:
+Next window prediction
 : Given that token $t$ was drawn from the distribution, the next latent window itself is constructed by a transformation $F:X^{w}\to X^{w}$, which advances the window as follows:
@@
 These two stages interact when they produce the LLM output, so it is important to understand the lineage of a given tokenization as being from a particular LLM.
-A.2. Computational resource requirements
+A.3 Computational resource requirements
 Applying our entire method (starting with the token embedding matrix and ending with the $p$-values for each of the three tests at every token) to each model took approximately 12 hours of wall clock time (per model) on an Intel Core i7-3820 with 32 GB of CPU RAM and no GPU running at 3.60GHz.
-Almost all of the runtime was spent computing the token distance matrix.
-The subsequent stages, including the three hypothesis tests, shown in Figure 2 run in only a few minutes on the same machine.
+Almost all of the runtime was spent computing the token pairwise distance matrix $D$ in Algorithm 1.
 The distance matrix calculation was performed using scipy.spatial.distance_matrix(), which produces a dense matrix, by computing the distance between all pairs of tokens in the latent space.
 Because of the relatively low runtime requirements, we did not feel the need to use sparsity-enhancing techniques to accelerate the distance matrix calculation.
-A.3. Interpretation of fiber bundles in terms of spatial variability
+The subsequent stages, including the three hypothesis tests run in only a few minutes on the same machine.
+A.4 Interpretation of fiber bundles in terms of spatial variability
 It is usual to describe measurements as exhibiting the combined effect of multiple spatial scales.
@@
 In what follows, we will call $S$ the base space and $V$ the fiber space.
-Figure 5.
-Our method applied to a fiber bundle in $\mathbb{R}^{2}$.
+Figure 5: Our method applied to a fiber bundle in $\mathbb{R}^{2}$.
 The vertical direction is the base space (long spatial scale), while the horizontal direction represents the fibers space (short spatial scale).
@@
 Figure 5 shows an example of this situation.
-It consists of a $1$-dimensional base space (the long spatial scale) and $1$-dimensional fibers (the short spatial scale), which in this case forms a narrow strip in the plane.
+It consists of a $1$-dimensional base space (the large radius neighborhood) and $1$-dimensional fibers (the small radius neighborhood), which in this case forms a narrow strip in the plane.
 Volumes (areas, in this case) of balls of small radius scale quadratically (slope $2$ in a log-log plot), but scale asymptotically linearly (slope $1$ in a log-log plot) for large radii.
@@
 In a fiber bundle, the long spatial scales are modeled by a space $S$, but all possible measurements are modeled by a function $p:E\to S$.
-The idea is that fibers $p^{-1}(b)$ are still cartesian products: pairs of both spatial scales, and these are all identical up to diffeomorphism.
+The idea is that fibers $p^{-1}(b)$ are still cartesian products: pairs of both spatial scales, and these are all identical up to smooth changes of coordinates.
 Our method relies upon a particular geometric property of fiber bundles: we can identify if the fibers are not all identical according to when the conclusion of Theorem 1 is violated.
-One possible way that a fiber bundle might arise is if the underlying noise model is homooscedastic, which is to say that the dimensionality of the noise near a token does not depend on that token.
-In this case, the fibers (short spatial scales) mostly correspond to noise, and the base (long spatial scales) correspond mostly to signal.
+One possible way that a fiber bundle might arise is if the underlying noise model is homoscedastic, which is to say that the dimension of the noise near a token does not depend on that token.
+In this case, the fibers (small radii) mostly correspond to noise, and the base (large radii) correspond mostly to signal.
 Nevertheless, a fiber bundle may arise with noiseless data as well; the presence of multiple spatial scales is what matters.
 On the other hand, if the underlying noise model is heteroscedastic—it depends on the token in question—then the fiber bundle model should be rejected.
-Figure 3(a) shows a situation that is not a fiber bundle, since there is a change in the dimension of the fiber.
-In the upper portion of the figure, the fiber dimension is 00 while in the lower portion the fiber dimension is $1$.
+Figure 2 shows a situation that is not a fiber bundle, since there is a change in the dimension of the fiber.
+In the left portion of the figure, the fiber dimension is 0 while in the right portion the fiber dimension is $1$.
 This is detectable by looking at the volume versus radius plots for two samples.
@@
 On the other hand, because the sample marked $\psi_{2}$ does not exhibit this violation, it is important to note that if a sample yields data consistent with Theorem 1, we cannot conclude that the space is a fiber bundle.
-A.4. Mathematical proof of Theorems 1 and 2
+A.5 Mathematical proof of Theorems 1 and 2
 This section contains mathematical justification for the fiber bundle hypothesis proposed earlier in the paper and the proof of Theorems 1 and 2.
 The central idea is the use of a special kind of fiber bundle, namely a fibered manifold with boundary.
-By the submersion theorem ([48]), if the Jacobian matrix of a fibered manifold $p:T\to S$ at every point has rank equal to the dimension of $S$, then the preimages $p^{-1}(\psi)\subseteq T$ of each point $\psi\in T$ are all diffeomorphic to each other.
+By the submersion theorem (Lee [2003]), if the Jacobian matrix of a fibered manifold $p:T\to S$ at every point has rank equal to the dimension of $S$, then the preimages $p^{-1}(\psi)\subseteq T$ of each point $\psi\in T$ are all diffeomorphic to each other.
 These preimages form the fibers discussed in the earlier sections of the paper.
@@
 The pushforward of a volume form is a contravariant functor, so this means that $e_{*}v$ is the volume form for a Riemannian metric on $e(T)$.
-Using this Riemannian metric on $e(T)$, then [49, Thm 3.1] implies that for every $\psi\in e(T)$, if $r\ll\tau$, then
-(2) $(e_{*}v)\left(B_{r}(\psi)\right)=O\left(r^{\dim T}\right).$
+Using this Riemannian metric on $e(T)$, then [Gray, 1974, Thm 3.1] implies that for every $\psi\in e(T)$, if $r\ll\tau$, then
+$(e_{*}v)\left(B_{r}(\psi)\right)=O\left(r^{\dim T}\right).$ (3)
 Since $T$ is compact, $S$ is also compact via the surjectivity of $p$.
@@
 Also by compactness of $S$, there is a minimum radius $r_{2}$ such that a ball of radius $r_{2}$ centered on a point $\psi\in e(T)$ contains a point outside of $e(T)$.
-Since $e$ is assumed to be an embedding, by the tubular neighborhood theorem ([48]), it must be that $r_{2}<\tau$. Define
+Since $e$ is assumed to be an embedding, by the tubular neighborhood theorem (Lee [2003]), it must be that $r_{2}<\tau$. Define
 $\rho(\psi):=\textrm{argmax }_{r}\left\{B_{r}(\psi)\subseteq e(T)\right\},$
@@
 We now address Theorem 2.
+Informally, we cannot get rid of all stochasticity during training (due to finite precision, stochastic gradient descent, etc.).
+This ensures that the transformers in practice are always in general position, hence generic.
+Given a context window of w1w-1 tokens, an LLM uses its learned parameters to predict the next ($w$-th) token.
+This gets added to the context window and the process continues.
+(See Section A.2 for details.) In this section, we provide a mathematical argument showing that the use of context is insufficient to resolve singularities present in the token subspace for a generic transformer in a persistent manner.
+Again, denote the token subspace by $T$, noting that Section 5.2 shows that $T$ is unlikely to be a manifold for typical LLMs.
+Let us begin with a lemma:
+Lemma 1.
+If $X$ is a manifold, $Y$ is an arbitrary topological space, and $X\times Y$ is a manifold, then $Y$ must also be a manifold.
 Proof.
-(of Theorem 2) We need to manipulate the inequalities in the hypothesis to show that they satisfy [50, Thm.
+Observe that $Y$ is homeomorphic to $(X\times Y)/X$, which is the quotient of two manifolds via the projection map, which is a surjective submersion. Hence $Y$ is a manifold. ∎
+Consequently, the contrapositive of Lemma 1 states that if $T$ is not a manifold but $T^{w-1}$ is a manifold, it follows that $T^{w}$, the space of sequences of tokens, cannot be a manifold.
+At best, the latent space for every other context window size can be a manifold.
+For every proposed context window length, a longer window will exhibit singularities.
+In short, lengthening the context window does not persistently resolve singularities in the token embedding.
+For instance, if we use the embedding interface provided by the LLM to embed a sequence of $w$ tokens, we will obtain a vector in the latent space $\mathbb{R}^{\ell}$.
+On the other hand, if we embed each of the tokens individually yet sequentially, we obtain a vector in $\mathbb{R}^{\ell\times w}$.
+The transformer uses the vector in $\mathbb{R}^{\ell\times w}$ to determine the vector in $\mathbb{R}^{\ell}$.
+Theorem 2 states that the latter ($\mathbb{R}^{\ell\times w}$) is crucial in understanding the output of the LLM.
+(See Section A.2 for further details.)
+Proof.
+(of Theorem 2) We need to manipulate the inequalities in the hypothesis to show that they satisfy [Robinson et al., 2025, Thm.
 1].
-Once the hypothesis is satisfied, [50, Thm.
+Once the hypothesis is satisfied, [Robinson et al., 2025, Thm.
 1] establishes that a generic set of transformers will embed $T^{w}$ into the output of the LLM.
-Notice that [50, Thm.
+Notice that [Robinson et al., 2025, Thm.
 1] considers the response of the LLM to a single token, though this can be easily extended by enlarging the bounding manifold dimension.
@@
 which can be easily rearranged to the form in the hypothesis. ∎
-A.5. Additional views for Figure 1
+A.6 Case study of some tokens in GPT2
+Figure 6: Application of the tests to real data consisting of the tokens #, ¢, and \$ with GPT2’s token input embedding.
+Notice that in the case of \$, the first slope change decreases the slope; the fiber bundle test fails if the slope increases.
+Figure 6 demonstrates our method on three tokens used by GPT2.
+The tokens \$ and # play an important role in many programming languages, whereas ¢ does not.
+(Note that these tokens were chosen for illustrative purposes.
+The $p$-value for rejecting the hypotheses is larger than $\alpha=10^{-3}$ used in Section 5.2, so these particular tokens do not appear in any of the tables in the following sections.) The token # does not show any rejections of our hypotheses.
+While this does not allow one to conclude that the vicinity of # is a manifold or fiber bundle, we can use Theorem 1 to estimate the dimension from the slope of the curve (approximately $53$, once logarithms of the values on both axes are taken).
+The token ¢ exhibits a rejection of the manifold hypothesis, but not the fiber bundle hypothesis.
+We can conclude that there is a irregularity at a distance of roughly $3.9$ units of ¢.
+The curve for \$ exhibits two slope changes at radii $3.8$ and $4.0$.
+The smaller radius slope change represents a violation of the manifold hypothesis, while the large radius slope change represents a violation of the fiber bundle hypothesis.
+Given the presence of multiple slope changes, it is necessary to select the parameters $v_{min}$ and $v_{max}$ in Algorithm 1 carefully.
+Ideally, one aims to capture at most one slope change between $v_{min}$ and $v_{max}$.
+This is not specifically required by Algorithm 1, since it uses a sliding window, multiple rejections may be detected.
+A.7 Case study of a neighborhood of ember in Mistral7B
+Figure 7: This figure shows the application of this paper’s manifold hypothesis test to a neighborhood of the token ember in Mistral7B.
+(The fiber bundle tests are not shown due to space considerations.) Since the latent space of Mistral7B has dimension $\ell=4096$, we have projected the coordinates of each token shown to $3$ using PCA.
+The inset shows all of the tokens in Mistral7B (rendered with t-SNE), with a crosshairs showing the location of the zoomed-in region.
+Each point shown in this figure is a token used by Mistral7B, colored based upon the $p$-value for the manifold hypothesis test.
+Darker colors correspond to lower $p$-values, which are less consistent with the hypothesis that the token’s neighborhood is a manifold with low curvature.
+Putative geometric structures are labeled for certain tokens.
+Figures 7 and 8 show the application of this paper’s manifold hypothesis test to a neighborhood of the token ember in Mistral7B.
+Since the latent space of Mistral7B is $4096$, we have projected the coordinates of each token shown to $3$ using PCA.
+The inset shows all of the tokens in Mistral7B (rendered with t-SNE), with a crosshairs showing the location of the zoomed-in region.
+Each point shown in Figure 7 is a token used by Mistral7B, colored based upon the $p$-value for the manifold hypothesis test.
+Darker colors correspond to lower $p$-values, which are less consistent with the hypothesis that the token’s neighborhood is a manifold with low curvature.
+Given that the we have established that manifold rejections occur in Section 5.2, we can attempt to understand possible causes.
+Intuitively, several non-manifold features are visibly present [^4] The coordinates, token text, $p$-values, and a video of the space with commentary are available in the ZIP file Supplement.
+Due to size requirements for the coordinates of the entire vocabularies for all four LLMs, we will upload the coordinates to a public repository.
+in Figure 7.
+The tokens “ember” and “P” appear to be both putative examples of cusp points, places where the space exhibits a sharp exposed point.
+Manifolds cannot have cusp points, so attempting to approximate a cusp with a manifold—for instance, with UMAP—necessitates high curvature.
+The tokens “:” and “ly” appear to be pinch points—like what are present in a chain of sausages—which again cannot be approximated by a manifold unless it has high curvature.
 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 6.
-Additional rotational views of the PCA of the neighborhood of ember in Mistral7B shown in Figure 1.
+Figure 8: Additional rotational views of the PCA of the neighborhood of ember in Mistral7B shown in Figure 7.
 Each point shown is a token.
 Colors are $p$-values for the manifold hypothesis, with darker colors corresponding to lower likelihood that that point’s neighborhood is a manifold with low curvature.
-See Fig. 6.
-A.6. Table of fiber bundle rejections in all four LLMs
-Table 3. Violations of the fiber bundle hypothesis
-Model Token Slope change $p$-value Comment
-GPT2 laughable Smaller radius $9\times 10^{-6}$ Must start a word
-GPT2 nuance Smaller radius $2\times 10^{-4}$ Must start a word
-GPT2 dt Smaller radius $2\times 10^{-4}$
-GPT2 Mesh Smaller radius $2\times 10^{-4}$
-GPT2 affect Smaller radius $3\times 10^{-4}$ Must start a word
-GPT2 Thankfully Smaller radius $3\times 10^{-4}$
-GPT2 swat Smaller radius $6\times 10^{-4}$ Must start a word
-GPT2 Malaysian Smaller radius $6\times 10^{-4}$ Must start a word
-GPT2 Palestinian Smaller radius $7\times 10^{-4}$ Must start a word
-GPT2 wins Smaller radius $8\times 10^{-4}$ Must start a word
-GPT2 hedon Smaller radius $9\times 10^{-4}$
-GPT2 donor Smaller radius $9\times 10^{-4}$ Must start a word
-GPT2 Xan Larger radius $3\times 10^{-8}$ Must start a word
-GPT2 aunder Larger radius $2\times 10^{-4}$
-GPT2 Dri Larger radius $2\times 10^{-4}$
-GPT2 ney Larger radius $3\times 10^{-4}$
-GPT2 rodu Larger radius $3\times 10^{-4}$
-GPT2 Insert Larger radius $4\times 10^{-4}$
-GPT2 Ying Larger radius $4\times 10^{-4}$ Must start a word
-Llemma7B pax Larger radius $3\times 10^{-4}$
-Mistral7B änge Smaller radius $8\times 10^{-4}$
-Mistral7B H0 Larger radius $5\times 10^{-4}$
-Mistral7B monitor Larger radius $8\times 10^{-5}$ Must start a word
+A.8 Table of fiber bundle rejections in all four LLMs
+Table 4: Violations of the fiber bundle hypothesis
+Model Token ID Token Slope change $p$-value Comment
+GPT2 47623 laughable Smaller radius $9\times 10^{-6}$ Must start a word
+GPT2 47128 nuance Smaller radius $2\times 10^{-4}$ Must start a word
+GPT2 28664 dt Smaller radius $2\times 10^{-4}$
+GPT2 37031 Mesh Smaller radius $2\times 10^{-4}$
+GPT2 2689 affect Smaller radius $3\times 10^{-4}$ Must start a word
+GPT2 48387 Thankfully Smaller radius $3\times 10^{-4}$
+GPT2 44284 swat Smaller radius $6\times 10^{-4}$ Must start a word
+GPT2 30031 Malaysian Smaller radius $6\times 10^{-4}$ Must start a word
+GPT2 8501 Palestinian Smaller radius $7\times 10^{-4}$ Must start a word
+GPT2 7864 wins Smaller radius $8\times 10^{-4}$ Must start a word
+GPT2 46086 hedon Smaller radius $9\times 10^{-4}$
+GPT2 17052 donor Smaller radius $9\times 10^{-4}$ Must start a word
+GPT2 47482 Xan Larger radius $3\times 10^{-8}$ Must start a word
+GPT2 21118 aunder Larger radius $2\times 10^{-4}$
+GPT2 20564 Dri Larger radius $2\times 10^{-4}$
+GPT2 1681 ney Larger radius $3\times 10^{-4}$
+GPT2 2076 rodu Larger radius $3\times 10^{-4}$
+GPT2 44402 Insert Larger radius $4\times 10^{-4}$
+GPT2 38833 Ying Larger radius $4\times 10^{-4}$ Must start a word
+Llemma7B 16352 pax Larger radius $3\times 10^{-4}$
+Mistral7B 25219 änge Smaller radius $8\times 10^{-4}$
+Mistral7B 9009 monitor Larger radius $8\times 10^{-5}$ Must start a word
+Mistral7B 4104 H0 Larger radius $5\times 10^{-4}$
 See Table 4.
-A.7. Tables of manifold rejections in all four LLMs
-Table 4. Rejections of the manifold hypothesis for GPT2 (1 of 2)
-Number Token $p$-value Comment
-1 Xan 0.0000000280 Must start a word
-2 chemist 0.0000000602
-3 fry 0.00000564 Must start a word
-4 laughable 0.00000807 Must start a word
-5 Retro 0.0000108 Must start a word
-6 laughable 0.0000156 Must start a word
-7 Thankfully 0.0000339
-8 intoler 0.0000514 Must start a word
-9 assumes 0.0000533 Must start a word
-10 handwritten 0.0000783 Must start a word
-11 MY 0.0000811 Must start a word
-12 nuance 0.000105 Must start a word
-13 playground 0.000110 Must start a word
-14 dt 0.000124
-15 aunder 0.000132
-16 Mesh 0.000135
-17 Disney 0.000139
-18 McCl 0.000155 Must start a word
-19 Dri 0.000196
-20 Destiny 0.000206 Must start a word
-21 ney 0.000206
-22 juvenile 0.000209 Must start a word
-23 advent 0.000211 Must start a word
-24 perpet 0.000213 Must start a word
-25 otally 0.000223
-26 affect 0.000226 Must start a word
-27 efficient 0.000238
-28 rodu 0.000242
-29 Kl 0.000260 Must start a word
-30 Thankfully 0.000278
-31 duct 0.000284
-32 frail 0.000292 Must start a word
-33 alle 0.000304 Must start a word
-34 Cha 0.000322 Must start a word
-35 OO 0.000333
-Table 5. Rejections of the manifold hypothesis for GPT2 (2 of 2)
-Number Token $p$-value Comment
-36 Insert 0.000336
-37 Ying 0.000357 Must start a word
-38 Georgia 0.000373
-39 flo 0.000384
-40 uddled 0.000413
-41 subsequ 0.000421 Must start a word
-42 lig 0.000439 Must start a word
-43 Environmental 0.000456 Must start a word
-44 Vladimir 0.000466 Must start a word
-45 Defence 0.000510 Must start a word
-46 swat 0.000558 Must start a word
-47 registering 0.000564 Must start a word
-48 Malaysian 0.000569 Must start a word
-49 trough 0.000575 Must start a word
-50 ears 0.000615 Must start a word
-51 Meh 0.000637 Must start a word
-52 Palestinian 0.000648 Must start a word
-53 reckless 0.000654 Must start a word
-54 Wax 0.000658 Must start a word
-55 Dist 0.000673 Must start a word
-56 Quad 0.000725 Must start a word
-57 relations 0.000749
-58 wins 0.000781 Must start a word
-59 KER 0.000785
-60 gist 0.000786 Must start a word
-61 hedon 0.000834
-62 MMA 0.000839 Must start a word
-63 donor 0.000865 Must start a word
-64 aleb 0.000873
-65 Asked 0.000873
-66 Isaiah 0.000940 Must start a word
-67 Survey 0.000947 Must start a word
-68 youngster 0.000987 Must start a word
-Table 6. Rejections of the manifold hypothesis for Llemma7B
-Number Token $p$-value Comment
-1 clojure 0.00000000434
-2 agu 0.000000846
-3 ¡Cyrillic¿ 0.000000938 Must start a word
-4 custom 0.00000163
-5 kW 0.00000496 Must start a word
+A.9 Tables of manifold rejections in all four LLMs
+Table 5: Rejections of the manifold hypothesis for GPT2 (1 of 2)
+Number Token ID Token $p$-value Comment
+1 47482 Xan 0.0000000280 Must start a word
+2 28899 chemist 0.0000000602
+3 32959 fry 0.00000564 Must start a word
+4 47623 laughable 0.00000807 Must start a word
+5 29884 Retro 0.0000108 Must start a word
+6 48387 Thankfully 0.0000339
+7 20619 intoler 0.0000514 Must start a word
+8 18533 assumes 0.0000533 Must start a word
+9 45916 handwritten 0.0000783 Must start a word
+10 17615 MY 0.0000811 Must start a word
+11 47128 nuance 0.000105 Must start a word
+12 24817 playground 0.000110 Must start a word
+13 28664 dt 0.000124
+14 21118 aunder 0.000132
+15 37031 Mesh 0.000135
+16 37338 Disney 0.000139
+17 23780 McCl 0.000155 Must start a word
+18 20564 Dri 0.000196
+19 17886 Destiny 0.000206 Must start a word
+20 1681 ney 0.000206
+21 21904 juvenile 0.000209 Must start a word
+22 19980 advent 0.000211 Must start a word
+23 8939 perpet 0.000213 Must start a word
+24 38908 otally 0.000223
+25 2689 affect 0.000226 Must start a word
+26 16814 efficient 0.000238
+27 2076 rodu 0.000242
+28 14770 Kl 0.000260 Must start a word
+29 2359 duct 0.000284
+30 41890 frail 0.000292 Must start a word
+31 28654 alle 0.000304 Must start a word
+32 20703 Cha 0.000322 Must start a word
+33 6684 $O$ 0.000333
+Table 6: Rejections of the manifold hypothesis for GPT2 (2 of 2)
+Number Token ID Token $p$-value Comment
+34 44402 Insert 0.000336
+35 38833 Ying 0.000357 Must start a word
+36 41072 Georgia 0.000373
+37 48679 flo 0.000384
+38 32745 uddled 0.000413
+39 6399 subsequ 0.000421 Must start a word
+40 26106 lig 0.000439 Must start a word
+41 13272 Environmental 0.000456 Must start a word
+42 14840 Vladimir 0.000466 Must start a word
+43 16721 Defence 0.000510 Must start a word
+44 44284 swat 0.000558 Must start a word
+45 28336 registering 0.000564 Must start a word
+46 30031 Malaysian 0.000569 Must start a word
+47 45047 trough 0.000575 Must start a word
+48 11368 ears 0.000615 Must start a word
+49 37070 Meh 0.000637 Must start a word
+50 8501 Palestinian 0.000648 Must start a word
+51 22053 reckless 0.000654 Must start a word
+52 42656 Wax 0.000658 Must start a word
+53 4307 Dist 0.000673 Must start a word
+54 20648 Quad 0.000725 Must start a word
+55 39468 relations 0.000749
+56 7864 wins 0.000781 Must start a word
+57 42839 KER 0.000785
+58 50057 gist 0.000786 Must start a word
+59 46086 hedon 0.000834
+60 19055 MMA 0.000839 Must start a word
+61 17052 donor 0.000865 Must start a word
+62 32100 aleb 0.000873
+63 18932 Asked 0.000873
+64 29756 Isaiah 0.000940 Must start a word
+65 13084 Survey 0.000947 Must start a word
+66 40944 youngster 0.000987 Must start a word
+Table 7: Rejections of the manifold hypothesis for Llemma7B
+Number Token ID Token $p$-value Comment
+1 24812 clojure 0.00000000434
+2 10617 agu 0.000000846
+3 24391 0.000000938 Cyrillic; Must start a word
+4 6341 custom 0.00000163
+5 26226 kW 0.00000496 Must start a word
+6 19579 }ˆ{- 0.00000833
+7 17485 porque 0.0000128 Must start a word
+8 23146 Leopold 0.0000135 Must start a word
+9 5240 Rem 0.0000232 Must start a word
+10 9317 Ern 0.0000265 Must start a word
+11 2466 though 0.0000540 Must start a word
+12 3413 č 0.0000619 Must start a word
+13 18659 endl 0.0000683 Must start a word
+14 24114 corrected 0.0000738 Must start a word
+15 3252 tw 0.0000788 Must start a word
+16 5908 ères 0.000115
+17 22385 grep 0.000151
+18 2924 kind 0.000202 Must start a word
+19 16352 pax 0.000243
+20 8602 Tri 0.000250 Must start a word
+21 18681 spielte 0.000258 Must start a word
+22 22977 Heaven 0.000262 Must start a word
+23 31946 0.000289 d with underline
+24 7949 0.000320 Cyrillic
+25 9160 Card 0.000335 Must start a word
+26 14406 Wall 0.000351 Must start a word
+27 29591 periodic 0.000359 Must start a word
+28 10236 ucht 0.000463
+29 18308 jeu 0.000553 Must start a word
+30 10029 slightly 0.000807 Must start a word
+31 15935 passion 0.000823 Must start a word
+32 5322 Charles 0.000833 Must start a word
+33 15354 victory 0.000906 Must start a word
+Table 8: Rejections of the manifold hypothesis for Mistral7B
+Number Token ID Token $p$-value Comment
+1 20639 CV 0.000000241 Must start a word
+2 1314 ember 0.00000199
+3 31684 0.00000224 Korean
+4 31567 ǐ 0.00000770
+5 8418 rise 0.0000283 Must start a word
+6 20508 poured 0.0000305 Must start a word
+7 17939 aimed 0.0000364 Must start a word
+8 21637 allocate 0.0000553 Must start a word
+9 13215 UINT 0.0000684
+10 9009 monitor 0.0000767 Must start a word
+11 25867 0.000105 Cyrillic
 […]
-7 porque 0.0000128 Must start a word
-8 Leopold 0.0000135 Must start a word
-9 Rem 0.0000232 Must start a word
-10 Ern 0.0000265 Must start a word
-11 though 0.0000540 Must start a word
-12 č 0.0000619 Must start a word
-13 endl 0.0000683 Must start a word
-14 corrected 0.0000738 Must start a word
-15 tw 0.0000788 Must start a word
-16 ères 0.000115
-17 grep 0.000151
-18 kind 0.000202 Must start a word
-19 pax 0.000243
-20 Tri 0.000250 Must start a word
-21 spielte 0.000258 Must start a word
-22 Heaven 0.000262 Must start a word
-23 d 0.000289
-24 ¡Cyrillic¿ 0.000320
-25 Card 0.000335 Must start a word
-26 Wall 0.000351 Must start a word
-27 periodic 0.000359 Must start a word
-28 ucht 0.000463
-29 jeu 0.000553 Must start a word
-30 slightly 0.000807 Must start a word
-31 passion 0.000823 Must start a word
-32 Charles 0.000833 Must start a word
-33 victory 0.000906 Must start a word
-Table 7. Rejections of the manifold hypothesis for Mistral7B
-Number Token $p$-value Comment
-1 CV 0.000000241 Must start a word
-2 ember 0.00000199
-3 ¡Korean character¿ 0.00000224
-4 ĭ 0.00000770
-5 rise 0.0000283 Must start a word
-6 poured 0.0000305 Must start a word
-7 aimed 0.0000364 Must start a word
-8 allocate 0.0000553 Must start a word
-9 UINT 0.0000684
-10 monitor 0.0000767 Must start a word
-11 ¡Cyrillic¿ 0.000105
-[…]
-13 ¡Chinese character¿ 0.000133
-14 Lat 0.000139 Must start a word
-15 Vel 0.000148
-16 challenges 0.000178 Must start a word
-17 ¡Chinese character¿ 0.000200
-18 Od 0.000222 Must start a word
-19 neys 0.000244
-20 selon 0.000244 Must start a word
-21 measures 0.000268 Must start a word
-22 assumes 0.000321 Must start a word
-23 Et 0.000338 Must start a word
-24 Rot 0.000438
-25 ¡Chinese character¿ 0.000538
-26 olutely 0.000540
-27 HO 0.000567
-28 Solutions 0.000598 Must start a word
-29 * 0.000658
-30 Condition 0.000665
-31 zt 0.000677
-32 iced 0.000696
-33 URL 0.000722
-34 SEO 0.000728 Must start a word
-35 weigh 0.000728 Must start a word
-36 änge 0.000765
-37 Vic 0.000777 Must start a word
-38 apparent 0.000780 Must start a word
-39 ¡Chinese character¿ 0.000896
-40 ¡0xE3¿ 0.000921
-Table 8. Rejections of the manifold hypothesis for Pythia6.9B (1 of 2)
-Number Token $p$-value Comment
-1 psychotic 0.000000161 Must start a word
-2 \\]( 0.00000147
-3 1990 0.00000231
-4 tiene 0.00000282 Must start a word
-5 ucky 0.0000134
-6 Hebrew 0.0000141 Must start a word
-7 mess 0.0000184
-8 reasonable 0.0000300 Must start a word
-9 collector 0.0000311 Must start a word
-10 embodiments 0.0000428 Must start a word
-11 Inhibition 0.0000593 Must start a word
-12 Hers 0.0000691 Must start a word
-13 odge 0.000120
-14 áĢIJ 0.000144
-15 ÏĢο 0.000165
-16 198 0.000170 Must start a word
-17 scare 0.000192 Must start a word
-18 ]{}\\_[ 0.000202
-19 LIMITED 0.000212 Must start a word
-20 hop 0.000232 Must start a word
-21 BIO 0.000246
-22 raison 0.000252 Must start a word
-23 inside 0.000254
-24 st 0.000261
-25 waving 0.000286 Must start a word
-26 Austria 0.000287 Must start a word
-27 ny 0.000313
-Table 9. Rejections of the manifold hypothesis for Pythia6.9B (2 of 2)
-Number Token $p$-value Comment
-28 bags 0.000326
-29 medium 0.000372
-30 incent 0.000378 Must start a word
-31 carboh 0.000407 Must start a word
-32 secret 0.000451 Must start a word
-33 ¡Unicode characters¿ 0.000468
-34 ÏĢε 0.000519 Must start a word
-35 120 0.000536
-36 wastes 0.000565 Must start a word
-37 ressor 0.000570
-38 Lip 0.000597 Must start a word
-39 preview 0.000650
-40 redund 0.000653 Must start a word
-41 ridic 0.000673 Must start a word
-42 Ramirez 0.000698 Must start a word
-43 omatic 0.000707
-44 Gloria 0.000733 Must start a word
-45 eton 0.000780
-46 chairman 0.000789 Must start a word
-47 Supreme 0.000859 Must start a word
-48 grazing 0.000890 Must start a word
-49 activation 0.000892 Must start a word
-50 imming 0.000914
-51 radiological 0.000931 Must start a word
-52 branc 0.000941 Must start a word
-53 \$\\—\\ 0.000942 Must start a word
-54 DEFAULT 0.000992
+13 31517 0.000133 Chinese
+14 7402 Lat 0.000139 Must start a word
+15 24351 Vel 0.000148
+16 10369 challenges 0.000178 Must start a word
+17 30085 0.000200 Chinese
+18 11424 Od 0.000222 Must start a word
+19 26983 neys 0.000244
+20 23851 selon 0.000244 Must start a word
+21 10582 measures 0.000268 Must start a word
+22 27460 assumes 0.000321 Must start a word
+23 13542 Et 0.000338 Must start a word
+24 13722 Rot 0.000438
+25 31097 0.000538 Chinese
+26 25244 olutely 0.000540
+27 4104 HO 0.000567
+28 24806 Solutions 0.000598 Must start a word
+29 29587 * 0.000658
+30 12862 Condition 0.000665
+31 3903 zt 0.000677
+32 3903 iced 0.000696
+33 5343 URL 0.000722
+34 21518 SEO 0.000728 Must start a word
+35 19929 weigh 0.000728 Must start a word
+36 19929 änge 0.000765
+37 19929 Vic 0.000777 Must start a word
+38 7566 apparent 0.000780 Must start a word
+39 31780 0.000896 Chinese
+40 230 <0xE3> 0.000921
+Table 9: Rejections of the manifold hypothesis for Pythia6.9B (1 of 2)
+Number Token ID Token $p$-value Comment
+1 49503 psychotic 0.000000161 Must start a word
+2 49503 \\]( 0.00000147
+3 13435 1990 0.00000231
+4 31918 tiene 0.00000282 Must start a word
+5 12202 ucky 0.0000134
+6 25260 Hebrew 0.0000141 Must start a word
+7 46631 mess 0.0000184
+8 5272 reasonable 0.0000300 Must start a word
+9 25389 collector 0.0000311 Must start a word
+10 24898 embodiments 0.0000428 Must start a word
+11 47769 Inhibition 0.0000593 Must start a word
+12 41264 Hers 0.0000691 Must start a word
+13 13676 odge 0.000120
+14 49669 0.000144 Turkish
+15 16287 0.000165 Turkish
+16 26397 198 0.000170 Must start a word
+17 30272 scare 0.000192 Must start a word
+18 41305 ]{}\\_[ 0.000202
+19 25666 LIMITED 0.000212 Must start a word
+20 5184 hop 0.000232 Must start a word
+21 39253 BIO 0.000246
+22 48089 raison 0.000252 Must start a word
+23 40084 inside 0.000254
+24 296 st 0.000261
+25 34475 waving 0.000286 Must start a word
+26 21753 Austria 0.000287 Must start a word
+27 5134 ny 0.000313
+Table 10: Rejections of the manifold hypothesis for Pythia6.9B (2 of 2)
+Number Token ID Token $p$-value Comment
+28 46728 bags 0.000326
+29 28672 medium 0.000372
+30 15210 incent 0.000378 Must start a word
+31 21742 carboh 0.000407 Must start a word
+32 4279 secret 0.000451 Must start a word
+33 50231 0.000468 misc. Unicode
+34 40146 0.000519 Turkish, Must start a word
+35 8193 120 0.000536
+36 47334 wastes 0.000565 Must start a word
+37 32232 ressor 0.000570
+38 22609 Lip 0.000597 Must start a word
+39 42150 preview 0.000650
+40 19886 redund 0.000653 Must start a word
+41 41962 ridic 0.000673 Must start a word
+42 48444 Ramirez 0.000698 Must start a word
+43 8977 omatic 0.000707
+44 46972 Gloria 0.000733 Must start a word
+45 18807 eton 0.000780
+46 16782 chairman 0.000789 Must start a word
+47 6413 Supreme 0.000859 Must start a word
+48 42351 grazing 0.000890 Must start a word
+49 5743 activation 0.000892 Must start a word
+50 34282 imming 0.000914
+51 46513 radiological 0.000931 Must start a word
+52 11363 branc 0.000941 Must start a word
+53 47183 $\|\\$ 0.000942 Must start a word
+54 25409 DEFAULT 0.000992
 See Tables 510.
+A.10 Limitations of theory
+Because our approach requires the distances between all pairs of tokens, to use the hypothesis tests presented in this paper, one needs to have access to the token embedding vectors.
+This is why our paper considers four open source models of moderate size – GPT2 (Radford et al.
+[2019]), Llemma7B (Azerbayev et al.
+[2024]), Mistral7B (Jiang et al.
+[2023]), and Pythia6.9B (Biderman et al.
+[2023]).
+While Robinson et al.
+[2025] does provide an approach to extract the token embeddings using systematic prompting, which would allow the application of our tests to proprietary models, their method requires exquisite control of the prompting and is computationally expensive.
+The reach itself is difficult to estimate accurately from sampled data (Berenfeld et al.
+[2022], Rawson and Robinson [2021], Aamari and Levrard [2019]).
+Our implementation tests the radius below an estimated reach, determined by an inspection of the radius-vs-volume plots of a random sample of tokens.
+If the true reach is smaller than our estimates, the curvature of the manifolds that fit the data well, found by any method, will be high.